A Complete Bibliography of Philosophy of Science: 2020–2029

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

01 April 2025 Version 1.11

Title word cross-reference

$105.00 [99]. $115.00 [313]. $119.99 [454]. $12.95 [314]. $129.99 [332]. $136 [399]. $159.99 [418]. $22.00 [312]. $24.99 [227]. $29.99 [45, 23]. $32.00 [226]. $34.95 [278]. $35.00 [240, 400, 241]. $39.95 [279]. $42.95 [46]. $42.95. [98]. $44.95 [208, 34]. $45 [420]. $45.00 [122]. $50.00 [295]. $57.00 [24]. $72.00 [206]. $74 [402]. $74.00 [333]. $75.00 [419]. $79.99 [242]. $88.00 [192]. $99 [401]. $99.00 [294, 207]. $99.99 [438, 518]. G [319, 165]. Λ [293]. Z [132].

//valuesinscience.com [295].

19 [487, 323].

2.1 [142]. 2020/2021 [382]. 2022 [459]. 2023 [439].

Abduction [199]. Ability [150]. Abnormality [357]. Absolutely [353].

Abstract [330]. Academia [125]. Academic [395]. access [295]. Account [318, 417, 432, 292, 62, 180, 122, 374, 120, 498, 468]. Accountable [469]. Accuracy [270, 223, 187]. Accuracy-First [187]. Accurate [239]. Action [163, 409]. Actionability [488]. Actual [299, 4, 250]. Actualism [330, 117]. Adam [418]. Adaptation [486, 31]. Adaptationism [359]. Adaptive [148]. Additivity [187, 424]. Address [459, 382]. Adequacy [30]. Adequacy-for-Purpose [30]. Adjacent [493]. Adrian [241]. Aesthetic [307]. Affects [125]. after [466, 51]. Against [376, 198, 475, 479, 363]. Agent [512]. Agent-Based [512]. Agents [350, 98]. Aggregation [507]. Agnosticism [464]. Ahistorical [388]. Aiming [86]. Aims [60, 59]. al [439]. Aldo [75, 494]. Algorithms [168]. also [206, 207, 208, 227]. Alternative [139, 179]. Alyssa [240]. American [146, 340, 415]. among [324, 94]. Amplification [258]. Analogue [251, 517]. Analysis [254, 378, 465, 256, 431, 396]. Analytic [253]. Analyzing [126]. Anarchy [63]. Andreas [518]. Animal [361, 443, 227]. Answer [216]. Answers [190]. Anthropocene [260]. Anti [173]. Anti-Humeans [173]. Antiscientism [340]. Apart [366]. Approach [142, 232, 386, 402]. Approaches [451, 436]. Approximate [131]. Approximately [42, 246]. Approximation [39]. Archaeology [203, 212]. Arguments [391, 118, 223]. Arnold [107]. Arrange [109]. Arranges [109]. Arrow [173]. Artful [109]. Artificial [38]. Ascription [357]. Ashtray [135, 275]. Asking [475]. Assessment [433]. Association [73, 382]. Astronomy [373]. Astrophysics [439, 84]. Asymbolia [426]. Asymmetry [322]. audiobook [227]. Austinian [380]. Autonomy [266]. Available [295, 206, 207, 208, 227]. Average [239]. Avoidance [134]. Avoidant [350]. Awareness [252].

B [516, 332]. B-S [516]. Babies [299]. Back [195, 211, 229, 244, 281, 298, 316, 336, 384, 404, 422, 441, 457, 504, 522]. Backward [269]. Baerdemaeker [439]. Bamboozled [502]. Bandits [427]. Bargaining [175, 286]. Barrett [192]. Base [317]. Base-Rate [317]. Based [86, 516, 512, 436]. Baseline [304]. Basic [22]. Batterman [402, 133]. Bayesian [183, 19, 343, 9, 20]. Be [272, 102, 246]. before [282]. Behavioral [231, 513, 91]. Behaviors [52]. Behind [419]. Being [355]. Beings [148]. Belief [507, 20, 436]. Beliefs [494]. Benefits [442, 520]. Best [412, 223, 351, 204, 473]. Better [102, 475, 304, 473, 239, 109]. Between [414, 303, 435, 491, 113]. Beyond [221, 259, 424, 45]. Bhopal [127]. Bias [356, 408, 125]. biases [356]. Binding [5]. Biological [92, 313, 69]. Biology [41, 312, 513, 249, 193]. Biomedical [487]. Biomimetic [453]. Birth [227]. Black [144, 516]. Body [402, 184]. Bohmian [28]. Bohr [420]. Bond [284]. Bone [241]. Bonferroni [502]. Book [206, 207, 226, 193, 136, 192, 418, 23, 99, 208, 98, 122, 24, 109, 227]. Books [156]. Borel [157]. Bound [216]. Boundary [103]. Box [85]. Boxes [144]. Boyd [439, 439, 180]. Break [369]. Breakdown [510]. Bridge [165]. Broadbent [110]. Broadly [126]. Brouwer [303]. Brown [295, 295].

Brownian [207]. Bueno [13]. Bugs [230].

Caitlin [419]. Calgary [454]. Calibration [29, 158]. Callender [120, 34]. Cambridge [279, 45, 312, 314, 400, 23, 122, 242, 438, 518, 241, 419]. Can [408, 102, 255, 160, 239, 343, 173, 246, 105, 109]. Cancellation [170]. Cancer [357]. Cardinality [434]. Carol [242]. Carolan [517]. Carrie [24]. Cartwright [314, 109, 314]. Case [190, 260, 331, 451, 319, 321, 125, 65, 323, 409]. Casey [399]. Categorical [274]. Categories [136]. Causal [139, 290, 142, 127, 312, 447, 352, 5, 173, 292, 358, 40, 461, 464, 405, 53, 54]. Causal-Role [292]. Causation [299, 72, 140, 428, 4, 247, 269, 54, 294, 294]. Cause [188]. Caused [127]. Centered [495]. Centric [237]. Certainty [97]. Certified [258]. Châtelet [482, 171]. Challenge [484, 476]. Challenges [139, 399]. Cham [439, 332, 517, 418]. Chance [349]. Change [125, 97]. Changes [268]. Changing [206]. Channeling [446]. Chaos [159]. Characterizing [310]. Characters [449]. Charles [312, 312]. Chemical [284]. Chemistry [460]. Chemostat [266]. Chicago [420]. Chike [99]. Choice [435, 290, 74, 443]. Choices [253]. Choosing [206]. Chronostratigraphic [260]. Clade [266]. Clades [214]. Claims [452]. Clarity [418]. Class [331, 68]. Classical [225, 366, 39, 19, 480, 340, 371]. Cleland [242]. Climate [159, 376, 64, 375, 486, 160, 23, 125, 237, 65, 50]. Closure [492]. cloth [420, 46, 122, 24, 34]. Clouds [160]. Coasters [73]. Code [289]. Coevolution [494]. Cognition [411, 343, 413, 200, 313, 313]. Cognitive [93, 449, 429, 432, 38]. Cohen [120]. Coherence [29, 73, 270, 145]. Coherentism [131, 499]. Colin [423]. Collaboration [362, 287]. Collapse [201]. Collective [265, 505, 262]. Collider [86]. Collins [227]. Comeasurability [118]. commentary [332]. Commercial [66, 468]. Commitments [330]. Community [259]. Commutativity [118]. Comparative [385]. Comparativism [44]. COMPASS [321]. Compatibility [508]. Complex [37, 8]. Complex-First [8]. Complexity [54, 338]. Component [246]. Components [141]. Composition [199, 216]. Compositional [58]. Compositionality [299, 4, 57]. Compressibility [90]. Computation [78, 328]. Computational [37, 506, 235, 289, 432]. Compute [349]. Computer [344, 102]. Concept [496, 104, 288, 508, 513, 358, 178, 122]. Concepts [488, 8]. Conceptual [466, 192, 338]. Concerns [445]. Condition [355]. Conditional [327, 55, 157, 20, 54]. conditionalization [448]. Conditioning [224]. Conditions [103]. Confidence [433]. Confident [116]. Confirmation [251, 114, 132, 408, 145, 113]. Conflict [113]. Conformity [196]. Connecting [303]. Consciousness [231, 227]. Consequentialists [300]. Conservation [256]. Considerations [342]. Consilience [29]. Constant [319, 2]. Constraint [299]. Constraints [460, 462, 461]. Construction [435, 31]. Contemporary [401]. Content [93, 307]. Contested [122]. Context [514, 480]. Contextual [445]. Contextuality [118]. Contingency [148]. Continuum [202, 323]. Control [163].

Convention [478, 56]. Conventional [303, 253]. Conventionalism [416]. Conventionality [481]. Convergence [362]. Conversations [193]. Cooperation [180]. Cooperative [346]. Coordinates [225]. Correcting [356]. Correspond [96]. Corrigendum [519]. Cosmological [83, 2]. Counterfactual [121]. Counterlegals [33]. Counterpossible [169]. Counts [185]. Cover [194, 195, 210, 211, 228, 229, 243, 244, 334, 280, 281, 297, 298, 315, 316, 335, 336, 383, 384, 403, 404, 421, 422, 440, 441, 456, 457, 503, 504, 521, 522]. COVID [487, 323]. COVID-19 [487, 323]. Craig [34]. Creativity [149]. Credal [113]. Credence [349, 436]. Credences [239]. Credit [68]. Creed [333]. Crisi [411]. Crisis [489, 255]. Critique [11, 180, 271]. Crossing [329]. Crowding [80]. Cultural [377, 16, 180, 166]. Cummins [232]. Currie [241]. Cycle [492].

D [454, 399]. Dappling [446]. Dark [305, 95, 485, 477]. Dark-Matter [305]. Darwinism [16]. Data [49, 420, 487, 237, 50, 162]. Data-Centric [237]. David [349]. Debate [221]. Debates [484]. Debiased [396]. Debiasing [166]. Deborah [45]. Deception [410]. Decision [164, 354, 9, 155]. Decision-Making [164]. Decisions [102, 116, 181]. Decoherent [406]. Decomposability [53]. Decoupling [283]. Deep [345, 182]. Default [151]. Defending [92, 475]. Defense [6, 10, 133, 307]. Defined [290]. Definition [141, 355]. Definitions [185]. Degree [20]. Demarcating [110]. Democratic [65, 347]. Democratizing [261]. Demographic [427]. Demons [235]. Demopoulos [279, 279]. Denial [125]. Denied [135]. Dependence [499]. Dependent [86]. Deprivation [347]. Deriving [172]. Descriptive [494]. Desiderata [94]. Design [148, 364, 5]. Details [127]. Detecting [463]. Detectors [163]. Determine [439]. Determinism [128]. Development [493, 288]. Developmental [446]. Dewey [123]. Dialogue [324]. Did [97]. Differences [303]. Difficulties [119]. Dilemma [428, 388, 285]. Dimension [176]. Dinosaurs [419]. Direct [231]. Directedness [496]. Disagree [208]. Disagreement [140]. Disagreements [326]. Disciplines [472]. Discounting [381]. Discourse [393]. Discovery [345]. Discrimination [310, 463]. Disease [414]. Disguise [447]. Disorders [230]. Dispositionalism [308]. Dissent [471, 125, 152]. Dissenting [509]. Distance [436]. Distance- [436]. Distention [222]. Distinction [445]. Distinguished [93]. Distortion [168]. Distribution [304]. Disunity [459, 442, 520]. Diverse [507]. Diversity [196, 427, 257, 150, 473]. Division [67]. Do [407, 219, 497]. Does [349, 96, 509, 412, 235, 186, 369]. Dogwhistles [389]. Domain [24, 168]. Dominik [517]. Donahue [419]. Don’t [242]. Down [369]. Downward [54]. Driven [95]. Drug [5]. Dual [31]. Duality [501, 81]. during [70]. Dutch [156]. Dynamic [142, 370, 413, 200]. Dynamical [370, 200]. Dynamics [356, 115, 77, 472]. Dysfunction [308].

e-book [206, 207, 208]. Easily [249]. ebook [517]. Ecological [176, 189, 292]. Ecology [513, 75]. Economics [321, 443]. Econophysics [129]. ed [332, 136]. edited [279, 439]. Edmond [177]. eds [418, 208, 401]. Effect [354, 378]. Effective [293, 510, 498]. Effects [177, 450]. Elaine [136]. Electromagnetism [81]. Elements [312]. Elizabeth [208]. Emanuele [401, 401]. Embrace [173]. Emerge [406]. Emergence [235, 213, 57]. Emergent [18]. Emergentism [201]. Emerging [428, 258]. Empathy [213]. Empirical [21, 139, 130, 28]. Empiricism [279]. Energy [318, 477, 184]. Engaged [126]. Engagement [347]. Engineer [387]. Engrams [77]. Ensembles [116]. Entity [324]. Entropy [96, 106, 17]. Environment [493]. Epidemics [197]. Epidemiological [144, 76]. Episodic [390]. Epistemic [435, 391, 346, 265, 513, 160, 61, 115, 257, 393, 374, 386, 300, 178, 473]. Epistemically [102, 343, 152, 328]. Epistemological [289]. Epistemology [453, 187, 149]. Epstein [11]. Equivalence [394, 82, 478, 17, 81]. Equivalent [143]. Eric [23]. Ernest [418]. Erosion [66]. Erotetic [462]. Erratum [334, 520]. Errol [135]. Error [32]. Essay [191, 45]. Essentialism [92, 508]. essentialist [122]. Ethic [75]. Ethnography [233, 212]. Ethos [258]. Etiological [89]. Evaluation [364, 30, 380]. Evaluative [494]. Evanston [278]. Evidence [11, 299, 509, 10, 326, 76, 183, 398, 205, 353, 275, 487]. Evidential [379, 515, 363]. Evolution [58, 175, 377, 400, 266, 472, 98]. Evolutionary [492, 320, 378, 513, 417, 180, 358, 213, 446, 495]. Excavation [373]. Exchangeability [224]. Expanded [164]. Expectation [55]. Expected [465]. Experience [167, 252]. Experimental [167, 299, 86, 203, 431]. Experimentation [465]. Experiments [299, 391, 95, 318, 487, 15]. Expertise [226, 487, 476]. Explaining [474, 313, 181]. Explanation [80, 344, 221, 144, 106, 278, 405, 204, 278]. Explanations [6, 447, 283, 462, 101, 234, 133, 358, 351, 392, 495]. Explanatory [352, 267, 198, 266, 145, 271, 515]. Exploration [501]. Exploratory [288, 465]. Exploring [378]. Exponential [381]. Externalist [483].

Face [294, 294]. Fair [238]. Fallacy [317]. Falls [366]. Family [306]. Fast [326]. Fathers [299]. Faulty [275]. Fauxrizon [397]. Favor [299]. Feedback [435]. Fictions [51]. Field [496, 293, 510, 382]. Fields [460]. Figdor [24]. Fight [418]. Filling [120]. Finding [73, 357]. Finite [378, 424]. First [187, 8]. Fischer [332]. Fitness [495]. Fitness-Centered [495]. FiveThirtyEight [255]. Fly [163]. Foes [257]. Foraging [360]. Forced [268]. foreword [295]. Form [307, 156]. Formal [289]. Formalization [260, 309]. Format [93]. Formation [288]. Formats [432, 227]. Forms [330]. Forward [341]. Forward-Looking [341]. Fossils [154, 178]. Foundational [205]. Foundations [192]. Four [99]. Fractured [48]. Framework [43, 126, 505]. French [13]. Frequentist [205]. Friedman [279]. Friends [257]. Front [14, 26, 100, 35, 47, 88, 112, 124, 137, 194, 210, 228, 243, 334, 280, 297, 315, 335, 383, 403, 421, 440, 456, 503, 521]. Fuller [331]. Function [514, 103, 189, 177, 450, 240]. Functional [427, 179]. Functionalism [325, 184]. Functionalist [374]. Functioning [75]. Functions [450, 357, 232, 75, 292]. Fundamental [402]. Funding [364, 304, 365]. Further [516]. Future [365]. Fuzzball [397].

G [45]. Gaia [309]. Gambit [483]. Game [377]. Game-Theoretic [377]. Games [286]. Gaps [474, 120]. Gas [127]. Gayon [193]. Gender [125]. General [172, 3, 128, 369]. Generalizations [431]. Generalized [16, 450, 55, 3]. Generating [239]. Generative [220]. Genericity [301]. Genome [73]. Genome-Wide [73]. Geoengineering [350]. Geologic [29]. George [207]. Get [45]. Glasgow [99, 227]. Glass [48]. Glennan [46]. Global [277]. Gloomy [73]. Goal [496]. Goals [61, 98]. Goblot [177]. Godfrey [227]. Godfrey-Smith [227]. Goldenberg [226]. Good [262]. GR [371]. Grammar [220]. Gravity [477]. Green [490]. Griceanism [120]. Gronda [123]. Grounding [327]. Group [408, 239, 180]. Groups [507]. Growth [411, 252]. Gualtiero [313, 313]. Guarantee [235]. Guide [241]. Guiding [413]. Guilt [417, 213]. Gunpoint [190].

H [332, 312]. Hadron [86]. Hangleiter [517]. Hannon [208]. hardback [517, 438]. Hardcover [295, 418, 294, 279, 206, 207, 333, 313, 240, 400, 208, 401, 242, 518, 402, 454, 227, 399, 241]. Harder [269]. Harm [509]. Harms [265]. Harvard [279]. Haslanger [99]. Hasty [431]. Hausdorff [3]. Health [362, 310, 414]. Hedge [452]. Heisenberg [332, 519]. Hempel [221]. Henderson [437]. Heng [439]. Heritability [72]. Hesitancy [226]. Heterogeneity [289]. Heuristic [486]. Heyting [303]. High [318]. Higher [326]. Higher-Order [326]. Historical [117, 278, 241, 373]. History [21, 352]. hit [247]. Hoboken [193]. Holism [413]. Hominin [180]. Howson [423]. https [295]. Human [32, 294, 208, 122, 91]. Hume [483]. Humean [273, 272]. Humeans [173]. Humility [147]. Hüttemann [518]. Hybridity [200]. Hydrology [202]. Hyperbole [275]. Hypotheses [185]. Hypothesis [412, 119, 186].

Ideal [295]. Idealization [272, 60]. Idealizations [202]. Idealized [38]. Ignore [261]. Illegitimate [445]. Imagination [295, 62, 63, 300]. Immunology [362]. Implausible [38]. Implementation [161]. Implications [160]. Implies [381]. Importance [127]. Impossibility [145]. Improve [408]. Inappropriate [152]. Incentives [395]. Incentivizing [151]. Incoherent [406]. incommensurability [448]. Inconsistent [507]. Incorrect [42]. Independence [9, 184, 54]. Independent [86, 141]. Independently [214]. Indexical [113]. Indicators [361]. Indigenous [476, 475]. Individual [331]. Individualism [215]. Individuality [492, 69, 495]. Individualized [513]. Individuating [449]. Individuation [359, 506]. Induction [437, 329, 454]. Inductive [301, 264]. Inequality [287]. Inertia [268]. Inextendibility [368]. Infant [411]. Inference [45, 301, 20, 501, 464, 204, 373]. Infinite [238, 270]. Influence [322, 364, 253]. Information [71, 347]. Informing [102]. Infrared [170]. Inheritance [16]. Inmaculada [12]. Innovation [365]. Institutional [364, 355, 468]. Institutions [469, 482]. Instrument [517]. Instruments [218]. Insufficiency [463]. Insufficient [151]. Intelligence [38]. Intelligibility [144]. Intemann [12]. Interacting [460]. Interactive [122]. Interests [66]. Internal [274]. Interpretations [376]. Interpreting [3]. Intertheoretic [510]. Interval [153]. Interventions [290]. Interviewed [193]. Intrinsic [92, 508]. introduction [332]. Invariance [108]. Invention [286]. Investigating [413]. Ioannidis [438]. Irreducibility [235]. Irrelevance [159]. Irreversible [247]. issue [194, 195, 210, 211, 228, 229, 243, 244, 280, 281, 297, 298, 315, 316, 335, 336, 383, 384, 403, 404, 421, 422, 440, 441, 456, 457, 503, 504, 521, 522]. ISTE [193]. ISTE/John [193].

J [226, 295, 400]. Jacques [517]. James [294, 294]. Jean [193]. Jeffers [99]. Jeffrey [333, 333, 192, 224]. Jill [225]. John [193, 454, 454]. Joshua [99]. Journals [395]. Judging [110]. Jurisprudence [168]. Just [406]. Justification [341, 106, 108, 452]. Justifying [425]. Juxtaposition [94].

Karim [517]. Kevin [439]. Kim [295, 400]. Kin [91]. Kinding [185]. Kinds [508, 89, 339]. Kinematics [224]. Kleinknecht [332]. Know [242]. Knowledge [203, 364, 475, 193]. Kochen [118]. Kolmogorov [157]. Kon [15]. Kon-Tiki [15]. Konrad [332, 332]. Kristen [12]. Kronfeldner [122]. Kuhn [275].

Labor [67]. Ladenness [49]. Land [75]. Landry [136]. Landscape [31]. Landscapes [427, 473]. Lange [133]. Language [58, 179, 400, 400]. Large [86]. Late [212]. Law [268, 168]. Laws [272, 33, 392, 109]. Learn [160]. Learning [390, 512, 345, 385, 408, 42, 55, 160, 125, 263, 8, 182, 162, 163, 264, 497, 181]. Learning-Theoretic [8]. Least [409]. Lectures [109]. Left [122]. Legitimacy [65]. Legitimate [445]. Legitimate/Illegitimate [445]. Leibniz [333]. Leopold [75, 494]. Lessons [449, 202, 510]. Levels [329, 325]. Lewens [208]. Lewis [349]. Life [193, 492, 214, 242, 461]. Like [230]. Limit [39, 397]. Limited [148]. Limiting [52]. Limits [293, 263, 261, 252, 166]. Lineage [180]. Lived [167]. Livengood [190]. Living [178]. Local [50]. Locating [511]. Logical [279, 156, 351]. London [193]. Looking [341, 48]. Looks [314, 269]. Loop [435]. Loose [235]. Lotteries [238, 365]. Lottery [304]. Low [6]. Low-Probability [6]. Luck [131]. Lukens [332].

M [332]. MA [122, 241, 419]. Machine [160, 263, 264, 497, 181]. Machines [160]. Made [203]. Maker [203]. Makes [231, 34]. Making [259, 164, 354, 116]. Mammalian [69]. Man [135]. Mandersian [394]. Manifolds [3]. Mantises [449]. Many [406, 70, 60, 107, 402]. Many-Body [402]. Map [377]. Map/Territory [377]. Marginalization [467]. Maria [122]. Mark [134]. Markers [393]. Martín [12]. Massimi [399]. Masthead [1]. Matarese [439]. Mate [74]. Material [203, 454]. materials [295]. Mathematical [344, 447, 101, 217]. Mathematics [130, 234, 171]. Mathematizing [409]. Matter [14, 26, 100, 35, 47, 88, 112, 124, 137, 334, 305, 95, 194, 195, 210, 211, 228, 229, 243, 244, 280, 281, 297, 298, 315, 316, 335, 336, 383, 384, 403, 404, 421, 422, 440, 441, 456, 457, 503, 504, 521, 522]. Matthew [295, 295]. Matthias [418]. Maya [226]. Mayo [45]. McCoy [399]. McDonough [333]. Meaningful [259]. Meaningless [246]. Means [396]. Measure [132, 354]. Measurement [71, 170, 374, 153, 499, 282, 207]. Measurements [28]. Measures [167, 29, 253, 231, 198, 271]. Measuring [154, 310, 282]. Mechanical [46, 168]. Mechanics [225, 451, 28, 192]. Mechanism [142]. Mechanisms [313, 283, 438, 438]. Mechanistic [267, 392]. Mechanistic-Explanatory [267]. Mediated [207]. Medicine [110, 191, 110]. Meeting [476]. Melo [12]. Melo-Martín [12]. Memetic [80]. Memoriam [423]. Memory [390, 185, 342, 275, 491]. Mental [230]. Merely [396]. Mesoscale [372]. Meta [356, 254, 253, 329, 256, 437]. Meta-Analysis [254, 256]. Meta-Analytic [253]. Meta-biases [356]. Meta-Induction [437, 329]. Metainduction [107]. Metalaws [273]. Metaphysics [240, 438, 518, 409]. Metarepresentational [234]. Metazoa [227]. Meteorology [50]. Method [95, 438, 166]. Method-Driven [95]. Methodological [117, 399]. Methodology [279, 391, 413, 217]. Methods [160, 239, 379, 107]. Michael [279]. Michela [399, 399]. Microaggressions [167, 474]. Middle [402]. Mills [439]. Mimicry [410]. Mind [267, 24, 184]. Mind-Body [184]. Minds [429, 227]. Minimal [133, 518]. Minority [471]. Miracle [333]. Miracles [317]. Misuses [205]. Mitigating [158]. Mitonuclear [508]. Mixed [379]. Mixed-Methods [379]. Mixtures [302]. Modality [394]. Model [86, 49, 516, 165, 356, 306, 164, 427, 321, 237, 133, 266, 30, 50, 116, 31, 395, 380]. Model-Data [50]. Model-Independent [86]. Model-Ladenness [49]. Modeler [109]. Modeling [372, 377, 429, 493, 237, 348, 323, 52, 94]. Models [93, 159, 376, 311, 389, 512, 407, 375, 102, 378, 318, 74, 51, 343, 323, 443, 38, 264, 497, 200, 94, 491]. Modern [279, 217]. Modified [477]. Modularity [179]. Molecular [207, 491]. Moral [320, 295]. Morris [135, 275]. Motion [207]. Motivating [323, 374]. Multi [321]. Multi-Model [321]. Multiple [41, 186, 40, 94]. Multiple-Models [94]. Multiscale [372].

N [332]. Nagel [418]. Nancy [314, 314, 109]. Nash [175]. Natural [176, 312, 508, 214, 339, 323]. Naturalism [414]. Naturalizing [32]. Naturalness [341]. Nature [32, 319, 122, 109, 208, 109]. Navigating [485]. Near [53]. Near-Decomposability [53]. Nearly [378]. Necessarily [236]. Necessary [33]. Need [412, 171]. Neither [153]. Nets [183]. Network [318]. Networks [115, 287]. Neuber [418]. Neumann [96, 17]. Neural [514, 512, 232, 77]. Neurocognitive [313, 313]. Neuroethology [372].

Neurons [38]. Neuroplasticity [186]. Neuroscience [288]. Neutral [378, 338]. Neutrality [381, 348]. Newton [97]. Ney [240]. Niche [435, 513, 31]. Niels [420]. NJ [193]. No [174, 317, 101, 334, 406]. No-Miracles [317]. No-Time [174]. NOAA [364]. Non [259, 3, 402, 250]. Non-actual [250]. Non-Fundamental [402]. Non-Hausdorff [3]. Non-Specialists [259]. Noncausal [462]. Nonexistent [282]. Nonneuronal [413]. Nontrivial [57]. nook [227]. Nor [153]. Nora [439]. Norm [258]. Normal [357]. Normality [357]. Normative [505]. Normativism [414]. North [225]. Northwestern [278]. Norton [454]. Notions [189]. Novel [398]. Nucleotide [72]. Null [74]. Number [7]. NY [294].

Objectivity [167, 291]. Observation [305, 220]. Observations [391, 84]. Occam [42]. Off [150]. Offs [94]. Okasha [98]. Old [398]. Oncology [488]. One [247]. One-hit [247]. Only [268]. Ontological [148, 215]. Ontology [339]. Opacity [345]. Opaque [264]. Open [259, 257, 295]. Opinion [270]. Optimality [333]. Optimist [241]. Order [326, 519, 332]. Ordered [22]. Ordinal [153]. Organic [460]. Organism [237]. Organisms [70, 164, 413]. Organizations [470]. Originated [214]. Origins [179, 180, 56]. Otavio [13]. Other [429]. Ought [219]. Outcome [474, 253]. Own [76]. Ownership [56]. Oxford [294, 206, 207, 333, 313, 192, 240, 46, 99, 208, 401, 98, 24, 402, 34, 34].

P [332, 517]. PAC [42]. Pain [426]. Paleoclimate [442, 520]. Pandemic [323, 468]. paper [45, 23]. paperback [419, 226, 312, 314, 227, 278]. Paradox [378, 157, 8, 204]. Parameter [501]. Participation [291]. Participatory [348]. Particle [510]. Particles [28]. Partition [108]. Past [412, 119]. Paternalist [396]. Path [499]. Path-Dependence [499]. Pattern [284]. Patterns [90]. Paul [278, 278]. Peace [45]. Peer [365, 395]. Pence [312, 312]. Perceived [360]. Perović [420]. Persistence [266]. Personal [248]. Perspectival [218]. Perspective [197, 164, 40]. Perspectives [401, 401]. Perspectivism [217, 399]. Pervasive [431]. Peter [227]. Petit [193]. Pettigrew [206]. Pharmaceutical [363]. phase [482]. Phenomenological [392]. Philosopher [314, 136]. Philosophers [219]. Philosophical [127, 99, 278]. Philosophy [21, 139, 333, 312, 123, 418, 191, 476, 233, 276, 431, 126, 248, 38, 415, 382, 439, 333, 46, 23, 382, 25, 209, 334, 296, 455, 111, 110]. Phylogenetic [443]. Phylogenetically [164]. Physical [103, 157, 328]. Physics [279, 333, 318, 510, 402, 217, 171, 169, 420, 134, 402, 279, 240, 225]. Physique [482]. Physiques [482]. Piccinini [313, 313]. Pieces [24]. Pittsburgh [226, 295]. Planer [400]. Planning [252]. Plants [449]. Pleistocene [212]. Pluralism [139, 363]. Pluralist [62, 122]. Points [236]. Poisoned [299]. Polarised [507]. Polarization [467]. Police [463]. Policy [365]. Political [65]. Polychrony [78]. Polymorphism [72]. Population [362, 197]. Populations [450, 378]. Position [28]. Positional [71].

Possibilist [376]. Possibilities [511]. Possible [493, 249]. Post [122]. Post-essentialist [122]. Postscript [224]. Power [352, 271, 201]. pp [294, 279, 206, 419, 207, 226, 332, 420, 333, 313, 192, 517, 45, 295, 312, 314, 418, 240, 46, 400, 23, 99, 208, 401, 98, 122, 242, 24, 438, 518, 402, 454, 227, 278, 34, 399, 241]. pp. [193]. Practice [154, 330, 473, 518]. Pragmatic [43, 127]. Pragmatism [146, 340]. Praying [449]. Precision [488]. Predicates [24]. Predicting [365]. Prediction [356, 285, 107]. Predictive [168]. Preface [337, 458, 138, 245]. Preferences [405]. Pregnancy [70, 69]. Preparing [419]. Presidential [459, 382]. Press [294, 279, 206, 207, 226, 420, 333, 313, 192, 45, 295, 312, 314, 240, 46, 400, 23, 99, 208, 401, 98, 122, 242, 24, 438, 518, 402, 454, 278, 34, 241, 419]. Pricean [309]. Pride [430]. Primitivism [479]. Principle [333, 106, 413, 409]. Principles [148, 341]. Privilege [374]. Probabilism [108, 223]. Probabilistic [44, 106, 198, 271, 433]. Probabilities [327, 331, 121, 222]. Probability [430, 6, 238, 108, 434, 157, 97]. Probably [42]. Problem [331, 154, 306, 140, 132, 329, 33, 68, 22, 398, 150, 2, 437, 201, 499]. Problems [205, 464]. Process [78]. Productive [63]. Progress [185, 425]. Projection [386]. Proof [104]. Proper [24]. Properties [18]. Proposals [139, 364]. Propositional [498]. Prospect [396]. Prospects [251, 73, 399, 399]. Proxies [442, 520]. PSA [459, 194, 195, 210, 211, 228, 229, 243, 244, 280, 281, 297, 298, 315, 316, 335, 336, 383, 384, 403, 404, 421, 422, 440, 441, 456, 457, 503, 504, 521, 522]. Psillos [438]. Psychiatric [355]. Psychiatry [355]. Psychological [24]. Psychologist [490]. Public [467, 226, 261, 262, 310, 452]. Purely [271]. Purpose [30]. Puzzle [157].

Qualitative [238]. Quality [198, 395]. Quanta [420]. Quantitative [479]. Quantities [481]. Quantum [129, 27, 387, 192, 517, 406, 460, 240, 105]. Quayshawn [99]. Quest [242]. Question [216]. Questions [475].

Race [99, 464]. Racial [310, 463]. Radiometric [29]. Raghav [207]. Ramseyfying [44]. Random [360, 204]. Randomize [188]. Randomness [360, 27]. Rate [317]. Rational [354, 343, 252]. Ratti [401]. Raven [204]. Razor [42]. Real [481, 284]. Real-Valued [481]. Realism [21, 484, 220, 324, 246]. Reality [207, 332, 325, 90, 519, 135, 225]. Realizability [186, 40]. Realization [41]. Really [129]. Reappraisal [177]. Reasoning [84, 321, 169]. Rebooting [38]. Reception [415]. Rechenberg [332]. Reciprocal [428]. Reckoning [202]. Reduction [451, 173]. Reductionism [325]. Referees [25, 209, 296, 455, 111]. Reference [331, 179, 68]. Reflections [516, 248]. Reflexivity [179]. Reformulations [143]. Reframing [221]. Regional [237]. Registering [471]. Regularity [238, 434]. Regulation [363]. Reinforcement [163]. Reintroducing [91]. Relationism [394]. Relations [510]. Relationship [377]. Relativism [166]. Relativity [366, 172, 3, 128, 415, 369, 53]. Relaxing [246]. Relevance [352, 113, 182]. Reliability [43]. Replication [151, 489, 254, 255, 411, 36, 67]. Reply [11, 398, 107]. Reports [471]. Represent [497]. Representation [274, 432, 120]. Representational [500]. Representations [232, 163, 53]. Representative [467]. Representing [250, 302]. Reproduction [69]. Requirement [11, 10]. Requirements [184]. Rescuing [215]. Research [104, 350, 379, 487, 237, 310]. Resemblance [161]. resides [491]. Resolving [484, 204]. Resource [304]. Response [331]. Responsibilities [505]. Responsibility [320, 265]. Responsible [469]. Results [253]. Rethinking [303, 466, 233]. Retraction [115]. Revealed [405]. Reversible [247]. Review [294, 519, 279, 439, 206, 419, 207, 226, 332, 420, 193, 333, 136, 313, 192, 517, 45, 295, 134, 110, 312, 314, 418, 240, 46, 191, 400, 23, 99, 208, 401, 98, 122, 242, 24, 438, 518, 365, 402, 454, 109, 227, 278, 34, 399, 241, 395, 225]. Revisited [416]. Revolutionary [303]. Revolver [190]. Rice [133]. Richard [206, 224]. Richerson [180]. Risk [76, 350, 264, 433]. Risk-Avoidant [350]. rivalry [491]. Road [348]. Robert [402, 402]. Roberto [123]. Robinson [295]. Robustness [305, 375, 83]. Rock [241]. Role [500, 172, 130, 355, 462, 234, 292, 117, 397]. Roles [513]. Roller [73]. Ronald [400, 400]. Roots [320]. Roth [278]. Routledge [399]. Ruin [241]. Ruined [299]. Rumscheidt [332].

S [516]. Safeguard [151]. Safeness [86]. Sally [99]. Samir [98]. Sampling [204]. Save [105]. Scaffold [358]. Scaffolded [493]. Scales [153]. Scenes [419]. scholarship [482]. Scholes [516]. Science [21, 470, 467, 466, 509, 344, 471, 469, 459, 259, 295, 326, 418, 289, 160, 256, 379, 68, 257, 487, 125, 237, 476, 348, 62, 307, 233, 401, 117, 475, 50, 386, 276, 22, 66, 126, 505, 438, 149, 261, 365, 262, 248, 342, 109, 473, 382, 445, 59, 468, 226, 295, 123, 314, 23, 401, 295, 314, 25, 209, 334, 296, 455, 111, 295]. Sciences [21, 103, 130, 461, 91, 241]. Scientific [21, 435, 391, 484, 6, 221, 154, 345, 259, 517, 341, 326, 183, 289, 258, 393, 234, 152, 323, 518, 150, 246, 120, 425, 217, 63, 452, 399]. Scientists [219, 261, 300]. Scope [365]. Scurvy [339]. Search [95]. Searches [86]. Searching [87, 242]. Secure [467]. Selected [177, 450]. Selection [176, 407, 306, 214, 5, 180, 266, 91, 464, 312, 312]. Selves [206]. Semantic [246]. Semiclassical [397]. Sensitivity [183]. Service [506]. Services [65]. Seth [207]. Sets [270, 222]. Setting [235]. Seven [21]. Severe [45]. Severity [205]. Sexes [407]. Sexual [407]. Shape [162]. Sharpening [450]. Shifts [370]. Shot [299]. Should [261]. Show [343]. Signal [400]. Signature [86]. Signature-Based [86]. Signatures [87]. Significance [372]. Similar [51]. Similarity [18]. Simple [204]. Simpler [102]. Simplicity [306]. Simulation [196, 517, 158, 102]. Simulations [439, 83, 84]. Single [72]. Siska [439]. Situations [515]. Sizes [354]. Sky [373]. Slobodan [420]. Small [7]. Smith [207, 227]. Social [512, 379, 461, 149, 472]. Society [466, 125]. softcover [418]. Software [230]. Soil [202]. Solomonoff [285]. Solution [8, 398]. Solving [255, 150]. Some [202, 248]. Sons [193]. Sophisticated [500]. South [277]. Space [511, 394, 466]. Spacetime [368, 480, 416]. Special [216, 34]. Specialists [259]. Species [508]. Specificity [5, 71]. Specker [118]. Speculation [203]. Spencer [99]. spin [371]. spin-2 [371]. Spite [378]. Sprenger [398]. Springer [439, 332, 517, 418]. Springs [444]. Stability [368]. Stabilization [515]. Stable [77]. Stack [239]. Stake [260]. Stances [484]. Stanley [295]. Stapleford [401]. Stars [439]. Start [154]. Start-Up [154]. State [216]. Statements [246]. Stathis [438]. Static [200]. Static-Dynamic [200]. Statistical [322, 311, 451, 352, 182, 501, 45, 45]. Statistics [45, 19, 463]. Status [366, 28]. Stavros [438, 438]. Sterelny [400]. Sterkenburg [107]. Steven [13]. Stochastic [9]. Strategic [460]. Strategies [286]. Strategy [86, 267]. Stratified [204]. String [172]. Structural [474, 147, 108, 287, 155]. Structure [82, 225, 274, 18, 183, 312, 351, 278]. Structuring [469]. Struggle [439]. Stuart [46, 46]. Studies [73]. Study [21, 260, 207, 196, 451, 443]. Studying [164]. Sub [306]. Sub-Family [306]. Subjective [361, 231, 108, 367]. Subjectivity [79]. Substantial [172]. Substrate [184]. Superiority [391]. supplementary [295]. Support [186]. Supposition [311]. Suppress [509]. Sustainability [348]. Sustaining [247]. Switzerland [332, 418]. Symbiosis [50]. Symbol [400]. Symmetric [327]. Symmetries [273, 370, 444]. Symmetry [172]. synaptic [491]. Synergistic [378]. Synthetic [249]. System [412, 7]. Systematic [431]. Systems [37, 41]. Sytsma [190].

T [366]. Take [268]. Tamas [418]. Targets [64, 179, 497, 250]. Taxonomy [49]. Teach [255, 495]. Teaching [276]. Technologies [350]. Technology [470, 401]. Teleofunction [506]. Teleosemantics [388]. Temperature [366]. Temporal [322, 381]. Territory [377]. Testing [45, 205, 45]. Textbooks [276]. Thébault [517]. Their [497]. Theoretic [377, 8]. Theoretical [219]. Theories [279, 435, 500, 183, 498]. Theory [86, 435, 207, 172, 177, 450, 293, 510, 434, 9, 242, 454, 396, 155, 371, 338]. Theory-Dependent [86]. Theory-Mediated [207]. Therapeutic [356]. There [101, 188, 491, 439]. Thermodynamic [96, 173, 17]. Thermodynamics [367, 105]. Think [32, 277, 219]. Thinking [494]. Thomas [401, 275]. Thumb [490]. Tiki [15]. Tim [208]. Time [29, 174, 154, 173, 105, 34]. Timescale [53]. Today [193]. Too [157]. Tool [288]. Topological [283]. Torsion [480]. Total [11, 10]. Totality [434]. Tracking [114]. Trade [150, 94]. Trade-Off [150]. Trade-Offs [94]. Tragedy [127]. Trait [359, 495]. Trait-Centered [495]. Transdisciplinary [476]. Transformative [252]. Transforming [487]. Transition [364]. Transitions [492, 495]. translated [332]. Transparency [37, 263, 113]. Traumatic [342]. Triangulation [448]. True [246, 120]. Trust [467, 226, 151, 196, 66, 262]. Trusting [342]. Trustworthiness [346]. Trustworthy [205]. Truth [61]. Tuboly [418]. Two [189, 451].

Unboundedly [107]. Uncertainties [64]. Uncertainty [158, 411, 515]. Unconscious [267]. Underdetermination [27]. Undermine [184]. Understanding [143, 61, 510, 411, 182, 264, 399, 59, 517, 399]. Unextended [236]. Unified [427]. Uniformitarianism [330]. Unit [260]. Unity [459]. Universal [242]. Universality [52]. Universe [85]. University [294, 279, 206, 207, 226, 420, 333, 313, 192, 45, 295, 312, 314, 240, 46, 23, 99, 208, 401, 98, 242, 24, 438, 518, 402, 454, 278, 34]. Unpredictability [235]. Unrestricted [390]. Use [76]. Useful [328]. Utility [350, 436]. Utility-Based [436].

Vaccine [226]. Validating [361, 85]. Validity [282]. Valuation [68]. Value [465, 348, 307, 263, 178]. Value-Neutrality [348]. Valued [481]. Values [511, 435, 470, 467, 364, 295, 64, 386, 445, 59, 468]. valuesinscience.com [295]. Variable [141, 290, 464]. Varieties [79, 237]. Variety [379]. Vassend [121]. Vera [439]. Verification [289]. Verisimilitude [121]. Versus [467, 19, 205]. VFI [466]. Victor [193]. View [220, 78, 65, 30]. Views [324, 97, 99]. viii [518]. Virtues [401, 219]. Vision [420]. volume [194, 195, 210, 211, 228, 229, 243, 244, 280, 281, 297, 298, 315, 316, 335, 336, 383, 384, 403, 404, 421, 422, 440, 441, 456, 457, 503, 504, 521, 522].

W [402]. War [226, 45]. Warp [168]. Wars [45]. Wave [240]. Wavefunction [387]. Way [86, 304, 402]. Welfare [361, 396]. Well [290, 355, 22]. Well-Being [355]. Well-Defined [290]. Well-Ordered [22]. Werner [519, 332]. Weyl [303, 172]. Where [491, 369]. Who [135]. Wide [73]. Wild [21]. Wiley [193]. William [279, 227, 279]. Wilson [134]. Winsberg [23]. Wishful [144]. Without [187, 362]. Woodward [294, 294]. Work [419, 191, 160]. Working [136]. Works [303]. World [240, 109, 459]. Worlds [406]. Worry [267]. Wylie [419, 419].

XI [517]. xx [193, 34].

York [294, 240, 46, 402, 399].

Zero [353].

References

Anonymous:2020:M

[1] Anonymous. Masthead. Philosophy of Science, 87(1):??, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schneider:2020:WPC

[2] Mike D. Schneider. What’s the problem with the cosmological constant?

Philosophy of Science, 87(1):1–20, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Luc:2020:INH

[3] Joanna Luc and Tomasz Placek. Interpreting non-Hausdorff (generalized) manifolds in general relativity. Philosophy of Science, 87(1):21–42, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Livengood:2020:ACC

[4] Jonathan Livengood and Justin Sytsma. Actual causation and compositionality. Philosophy of Science, 87(1):43–69, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). See response [190].

Lean:2020:BSC

[5] Oliver M. Lean. Binding specificity and causal selection in drug design. Philosophy of Science, 87(1):70–90, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Clatterbuck:2020:DLP

[6] Hayley Clatterbuck. A defense of low-probability scientific explanations. Philosophy of Science, 87(1):91–112, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Margolis:2020:SNS

[7] Eric Margolis. The small number system. Philosophy of Science, 87(1): 113–134, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Poth:2020:LCL

[8] Nina L. Poth and Peter Brössel. Learning concepts: a learning-theoretic solution to the complex-first paradox. Philosophy of Science, 87(1):135– 151, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mongin:2020:BDT

[9] Philippe Mongin. Bayesian decision theory and stochastic independence. Philosophy of Science, 87(1):152–178, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Draper:2020:DRT

[10] Paul Draper. In defense of the requirement of total evidence. Philosophy of Science, 87(1):179–190, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Barrett:2020:RTE

[11] Martin Barrett and Elliott Sober. The requirement of total evidence: a reply to Epstein’s critique. Philosophy of Science, 87(1):191–203, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dupre:2020:IMM

[12] John Dupré. Inmaculada de Melo-Martín and kristen intemann. Philosophy of Science, 87(1):204–207, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kasirzadeh:2020:OBS

[13] Atoosa Kasirzadeh and James Robert Brown. Otavio Bueno and Steven French. Philosophy of Science, 87(1):207–211, January 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2020:FMa

[14] Anonymous. Front matter. Philosophy of Science, 87(2):??, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Novick:2020:KTE

[15] Aaron Novick, Adrian M. Currie, Eden W. McQueen, and Nathan L. Brouwer. Kon-Tiki experiments. Philosophy of Science, 87(2):213–236, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Feldbacher-Escamilla:2020:CIG

[16] Christian J. Feldbacher-Escamilla and Karim Baraghith. Cultural inheritance in generalized Darwinism. Philosophy of Science, 87(2):237–261, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Prunkl:2020:ENT

[17] Carina E. A. Prunkl. On the equivalence of von Neumann and thermodynamic entropy. Philosophy of Science, 87(2):262–280, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fletcher:2020:SSE

[18] Samuel C. Fletcher. Similarity structure and emergent properties. Philosophy of Science, 87(2):281–301, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Johannesson:2020:CVB

[19] Eric Johannesson. Classical versus Bayesian statistics. Philosophy of Science, 87(2):302–318, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Sprenger:2020:CDB

[20] Jan Sprenger. Conditional degree of belief and Bayesian inference. Philosophy of Science, 87(2):319–335, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Beebe:2020:SRW

[21] James R. Beebe and Finnur Dellsén. Scientific realism in the wild: an empirical study of seven sciences and history and philosophy of science. Philosophy of Science, 87(2):336–364, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Philippi:2020:WOS

[22] Cristian Larroulet Philippi. Well-ordered science’s basic problem. Philosophy of Science, 87(2):365–375, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lam:2020:BRE

[23] Vincent Lam. Book review: Eric Winsberg, Philosophy and Climate Science. Cambridge: Cambridge University Press (2018), 270 pp., $29.99 (paper). Philosophy of Science, 87(2):376–379, April 2020. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pence:2020:BRC

[24] David Evan Pence. Book review: Carrie Figdor, Pieces of Mind: The Proper Domain of Psychological Predicates. Oxford: Oxford University Press (2018), 240 pp., $57.00 (cloth). Philosophy of Science, 87(2):379– 382, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2020:RPS

[25] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 87(2):383–389, April 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2020:FMb

[26] Anonymous. Front matter. Philosophy of Science, 87(3):??, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Barrett:2020:QRU

[27] Jeffrey A. Barrett and Simon M. Huttegger. Quantum randomness and underdetermination. Philosophy of Science, 87(3):391–408, July 2020. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lazarovici:2020:PME

[28] Dustin Lazarovici. Position measurements and the empirical status of particles in Bohmian mechanics. Philosophy of Science, 87(3):409–424, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bokulich:2020:CCC

[29] Alisa Bokulich. Calibration, coherence, and consilience in radiometric measures of geologic time. Philosophy of Science, 87(3):425–456, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Parker:2020:MEA

[30] Wendy S. Parker. Model evaluation: an adequacy-for-purpose view. Philosophy of Science, 87(3):457–477, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Tanaka:2020:DLM

[31] Mark M. Tanaka, Peter Godfrey-Smith, and Benjamin Kerr. The dual landscape model of adaptation and niche construction. Philosophy of Science, 87(3):478–498, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Allchin:2020:HWT

[32] Douglas Allchin and Alexander J. Werth. How we think about human nature: The naturalizing error. Philosophy of Science, 87(3):499–517, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kimpton-Nye:2020:NLP

[33] Samuel Kimpton-Nye. Necessary laws and the problem of counterlegals. Philosophy of Science, 87(3):518–535, July 2020. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weatherall:2020:RCC

[34] James Owen Weatherall. Review of Craig Callender’s What Makes Time Special? Oxford: Oxford University Press (2017), xx + 343 pp., $44.95 (cloth). Philosophy of Science, 87(3):536–544, July 2020. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2021:FMa

[35] Anonymous. Front matter. Philosophy of Science, 87(4):??, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Machery:2021:WR

[36] Edouard Machery. What is a replication? Philosophy of Science, 87 (4):545–567, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Creel:2021:TCC

[37] Kathleen A. Creel. Transparency in complex computational systems. Philosophy of Science, 87(4):568–589, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stinson:2021:IAN

[38] Catherine Stinson. From implausible artificial neurons to idealized cognitive models: Rebooting philosophy of artificial intelligence. Philosophy of Science, 87(4):590–611, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Feintzeig:2021:CLA

[39] Benjamin H. Feintzeig. The classical limit as an approximation. Philosophy of Science, 87(4):612–639, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ross:2021:MRC

[40] Lauren N. Ross. Multiple realizability from a causal perspective. Philosophy of Science, 87(4):640–662, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fang:2021:MRS

[41] Wei Fang. Multiple realization in systems biology. Philosophy of Science, 87(4):663–684, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Herrmann:2021:PLO

[42] Daniel A. Herrmann. PAC learning and Occam’s Razor: Probably approximately incorrect. Philosophy of Science, 87(4):685–703, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Davis:2021:FPR

[43] Isaac Davis. A framework for pragmatic reliability. Philosophy of Science, 87(4):704–726, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott:2021:RPC

[44] Edward Elliott. ‘Ramseyfying’ probabilistic comparativism. Philosophy of Science, 87(4):727–754, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fletcher:2021:WPE

[45] Samuel C. Fletcher. Of war or peace? Essay review of Statistical Inference as Severe Testing. Deborah G. Mayo, Statistical Inference as Severe Testing: How to Get beyond the Statistics Wars. Cambridge: Cambridge University Press (2018), 500 pp., $29.99 (paper). Philosophy of Science, 87(4):755–762, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Huneman:2021:RSG

[46] Philippe Huneman. Review of Stuart Glennan, The New Mechanical Philosophy. Stuart Glennan, The New Mechanical Philosophy. New York: Oxford University Press (2017), 256 pp., $42.95 (cloth). Philosophy of Science, 87(4):763–769, October 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2021:FMb

[47] Anonymous. Front matter. Philosophy of Science, 87(5):??, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mitchell:2021:TFL

[48] Sandra D. Mitchell. Through the fractured looking glass. Philosophy of Science, 87(5):771–792, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bokulich:2021:TTM

[49] Alisa Bokulich. Towards a taxonomy of the model-ladenness of data. Philosophy of Science, 87(5):793–806, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Parker:2021:LMD

[50] Wendy Parker. Local model-data symbiosis in meteorology and climate science. Philosophy of Science, 87(5):807–818, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Levy:2021:MFS

[51] Arnon Levy. Models and fictions: Not so similar after all? Philosophy of Science, 87(5):819–828, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rice:2021:UML

[52] Collin Rice. Universality and modeling limiting behaviors. Philosophy of Science, 87(5):829–840, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weinberger:2021:NDT

[53] Naftali Weinberger. Near-decomposability and the timescale relativity of causal representations. Philosophy of Science, 87(5):841–856, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Woodward:2021:CCC

[54] James Woodward. Causal complexity, conditional independence, and downward causation. Philosophy of Science, 87(5):857–867, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Huttegger:2021:GLC

[55] Simon M. Huttegger and Michael Nielsen. Generalized learning and conditional expectation. Philosophy of Science, 87(5):868–883, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Smead:2021:COO

[56] Rory Smead and Patrick Forber. Convention and the origins of ownership. Philosophy of Science, 87(5):884–896, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Steinert-Threlkeld:2021:TEN

[57] Shane Steinert-Threlkeld. Toward the emergence of nontrivial compositionality. Philosophy of Science, 87(5):897–909, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Barrett:2021:ECL

[58] Jeffrey A. Barrett, Calvin Cochran, and Brian Skyrms. On the evolution of compositional language. Philosophy of Science, 87(5):910–920, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

deRegt:2021:UVA

[59] Henk W. de Regt. Understanding, values, and the aims of science. Philosophy of Science, 87(5):921–932, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Potochnik:2021:IMA

[60] Angela Potochnik. Idealization and many aims. Philosophy of Science, 87 (5):933–943, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Khalifa:2021:UTE

[61] Kareem Khalifa. Understanding, truth, and epistemic goals. Philosophy of Science, 87(5):944–956, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Murphy:2021:TPA

[62] Alice Murphy. Toward a pluralist account of the imagination in science. Philosophy of Science, 87(5):957–967, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Stuart:2021:PAS

[63] Michael T. Stuart. The productive anarchy of scientific imagination. Philosophy of Science, 87(5):968–978, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Frisch:2021:UVC

[64] Mathias Frisch. Uncertainties, values, and climate targets. Philosophy of Science, 87(5):979–990, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lusk:2021:PLD

[65] Greg Lusk. Political legitimacy in the democratic view: The case of climate services. Philosophy of Science, 87(5):991–1002, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pinto:2021:CIE

[66] Manuela Fernández Pinto. Commercial interests and the erosion of trust in science. Philosophy of Science, 87(5):1003–1013, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Romero:2021:DRL

[67] Felipe Romero. The division of replication labor. Philosophy of Science, 87(5):1014–1025, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lee:2021:RCP

[68] Carole J. Lee. The reference class problem for credit valuation in science. Philosophy of Science, 87(5):1026–1036, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kingma:2021:BIP

[69] Elselijn Kingma. Biological individuality, pregnancy, and (mammalian) reproduction. Philosophy of Science, 87(5):1037–1048, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Grose:2021:HMO

[70] Jonathan Grose. How many organisms during a pregnancy? Philosophy of Science, 87(5):1049–1060, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Love:2021:PIM

[71] Alan C. Love. Positional information and the measurement of specificity. Philosophy of Science, 87(5):1061–1072, December 2021. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bourrat:2021:CSN

[72] Pierrick Bourrat. Causation and single nucleotide polymorphism heritability. Philosophy of Science, 87(5):1073–1083, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Craver:2021:GPR

[73] Carl F. Craver, Mikhail Dozmorov, Mark Reimers, and Kenneth S. Kendler. Gloomy prospects and roller coasters: Finding coherence in genome-wide association studies. Philosophy of Science, 87(5):1084–1095, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kovaka:2021:MCN

[74] Karen Kovaka. Mate choice and null models. Philosophy of Science, 87(5): 1096–1106, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Millstein:2021:FFA

[75] Roberta L. Millstein. Functions and functioning in Aldo Leopold’s land ethic and in ecology. Philosophy of Science, 87(5):1107–1118, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fuller:2021:EEU

[76] Jonathan Fuller. Epidemiological evidence: Use at your ‘own risk’? Philosophy of Science, 87(5):1119–1129, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Robins:2021:SEN

[77] Sarah K. Robins. Stable engrams and neural dynamics. Philosophy of Science, 87(5):1130–1139, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Klein:2021:PPV

[78] Colin Klein. Polychrony and the process view of computation. Philosophy of Science, 87(5):1140–1149, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Godfrey-Smith:2021:VS

[79] Peter Godfrey-Smith. Varieties of subjectivity. Philosophy of Science, 87(5):1150–1159, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Cao:2021:CME

[80] Rosa Cao. Crowding out memetic explanation. Philosophy of Science, 87(5):1160–1171, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weatherall:2021:EDE

[81] James Owen Weatherall. Equivalence and duality in electromagnetism. Philosophy of Science, 87(5):1172–1183, December 2021. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Barrett:2021:SE

[82] Thomas William Barrett. Structure and equivalence. Philosophy of Science, 87(5):1184–1196, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gueguen:2021:RCS

[83] Marie Gueguen. On robustness in cosmological simulations. Philosophy of Science, 87(5):1197–1208, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Jacquart:2021:OSR

[84] Melissa Jacquart. Observations, simulations, and reasoning in astrophysics. Philosophy of Science, 87(5):1209–1220, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Smeenk:2021:VUB

[85] Chris Smeenk and Sarah C. Gallagher. Validating the universe in a box. Philosophy of Science, 87(5):1221–1233, December 2021. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Beauchemin:2021:SBM

[86] Pierre-Hugues Beauchemin. Signature-based model-independent searches at the Large Hadron Collider: an experimental strategy aiming at safeness in a theory-dependent way. Philosophy of Science, 87(5):1234–1245, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Mattig:2021:SS

[87] Peter Mättig and Michael Stöltzner. Searching for signatures. Philosophy of Science, 87(5):1246–1256, December 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2021:FMc

[88] Anonymous. Front matter. Philosophy of Science, 88(1):??, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Khalidi:2021:EK

[89] Muhammad Ali Khalidi. Etiological kinds. Philosophy of Science, 88(1):1– 21, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Millhouse:2021:CRP

[90] Tyler Millhouse. Compressibility and the reality of patterns. Philosophy of Science, 88(1):22–43, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rubin:2021:RKS

[91] Hannah Rubin. Reintroducing kin selection to the human behavioral sciences. Philosophy of Science, 88(1):44–66, January 2021. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Devitt:2021:DIB

[92] Michael Devitt. Defending intrinsic biological essentialism. Philosophy of Science, 88(1):67–82, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Butlin:2021:CMD

[93] Patrick Butlin. Cognitive models are distinguished by content, not format. Philosophy of Science, 88(1):83–102, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Yoshida:2021:MMJ

[94] Yoshinari Yoshida. Multiple-models juxtaposition and trade-offs among modeling desiderata. Philosophy of Science, 88(1):103–123, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

DeBaerdemaeker:2021:MDE

[95] Siska De Baerdemaeker. Method-driven experiments and the search for dark matter. Philosophy of Science, 88(1):124–144, January 2021. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Chua:2021:DNE

[96] Eugene Y. S. Chua. Does von Neumann entropy correspond to thermodynamic entropy? Philosophy of Science, 88(1):145–168, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Shapiro:2021:NDC

[97] Alan E. Shapiro. Newton did change his views from certainty to probability. Philosophy of Science, 88(1):169–180, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

OConnor:2021:BRS

[98] Cailin O’Connor. Book review: Samir Okasha. Agents and Goals in Evolution. Oxford: Oxford University Press (2018), 272 pp., $42.95. Philosophy of Science, 88(1):181–184, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ludwig:2021:BRJ

[99] David Ludwig. Book review: Joshua Glasgow, Sally Haslanger, Chike Jeffers, and Quayshawn Spencer. What Is Race? Four Philosophical Views. Oxford: Oxford University Press (2019), 296 pp., $105.00. Philosophy of Science, 88(1):184–188, January 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2021:FM

[100] Anonymous. Front matter. Philosophy of Science, 88(2):??, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kuorikoski:2021:TNM

[101] Jaakko Kuorikoski. There are no mathematical explanations. Philosophy of Science, 88(2):189–212, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Helgeson:2021:WSC

[102] Casey Helgeson, Vivek Srikrishnan, Klaus Keller, and Nancy Tuana. Why simpler computer simulation models can be epistemically better for informing decisions. Philosophy of Science, 88(2):213–233, ???? 2021. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bursten:2021:FBC

[103] Julia R. S. Bursten. The function of boundary conditions in the physical sciences. Philosophy of Science, 88(2):234–257, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Elliott:2021:PCR

[104] Steve Elliott. Proof of concept research. Philosophy of Science, 88(2):258– 280, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Swanson:2021:CQT

[105] Noel Swanson. Can quantum thermodynamics save time? Philosophy of Science, 88(2):281–302, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Jodoin:2021:JPE

[106] Laurent Jodoin. A justification of the probabilistic explanation of the entropy principle. Philosophy of Science, 88(2):303–319, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schurz:2021:MUM

[107] Gerhard Schurz. Metainduction over unboundedly many prediction methods: A reply to Arnold and Sterkenburg. Philosophy of Science, 88(2): 320–340, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Leitgeb:2021:SJP

[108] Hannes Leitgeb. A structural justification of probabilism: From partition invariance to subjective probability. Philosophy of Science, 88(2):341– 365, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Veit:2021:BRN

[109] Walter Veit. Book review: Nancy Cartwright. Nature, the Artful Modeler: Lectures on Laws, Science, How Nature Arranges the World and How We Can Arrange It Better. Philosophy of Science, 88(2):366–369, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fuller:2021:DJM

[110] Jonathan Fuller. Demarcating and judging medicine: Review of Broadbent’s “Philosophy of Medicine”. Philosophy of Science, 88(2):370–376, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bramson:2021:RPS

[111] Aaron Bramson, Patrick Grim, Daniel J. Singer, William J. Berger, Graham Sack, Steven Fisher, Carissa Flocken, Bennett Holman, Wendy S.

Parker, James R. Beebe, Finnur Dellsén, Peter Machamer, Lindley Darden, Carl F. Craver, Carl G. Hempel, Paul Oppenheim, and Christopher Boorse. Referees for Philosophy of Science. Philosophy of Science, 88(2): 377–383, ???? 2021. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2022:FMa

[112] Anonymous. Front matter. Philosophy of Science, 88(3):??, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pust:2022:CBI

[113] Joel Pust. A conflict between indexical credal transparency and relevance confirmation. Philosophy of Science, 88(3):385–397, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Douven:2022:TC

[114] Igor Douven. Tracking confirmation. Philosophy of Science, 88(3):398– 414, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

LaCroix:2022:DRE

[115] Travis LaCroix, Anders Geil, and Cailin O’Connor. The dynamics of retraction in epistemic networks. Philosophy of Science, 88(3):415–438, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Roussos:2022:MCD

[116] Joe Roussos, Richard Bradley, and Roman Frigg. Making confident decisions with model ensembles. Philosophy of Science, 88(3):439–460, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Page:2022:RHS

[117] Meghan D. Page. The role of historical science in methodological actualism. Philosophy of Science, 88(3):461–482, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hofer-Szabo:2022:CCC

[118] Gábor Hofer-Szabó. Commutativity, comeasurability, and contextuality in the Kochen–Specker arguments. Philosophy of Science, 88(3):483–510, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Gryb:2022:NDP

[119] Sean Gryb. New difficulties for the past hypothesis. Philosophy of Science, 88(3):511–532, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Ruyant:2022:TGF

[120] Quentin Ruyant. True Griceanism: Filling the gaps in Callender and Cohen’s account of scientific representation. Philosophy of Science, 88(3): 533–553, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Niiniluoto:2022:VVC

[121] Ilkka Niiniluoto. Vassend on verisimilitude and counterfactual probabilities. Philosophy of Science, 88(3):554–561, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Odenbaugh:2022:BRM

[122] Jay Odenbaugh. Book review: Maria Kronfeldner. What’s Left of Human Nature? A Post-essentialist, Pluralist, and Interactive Account of a Contested Concept, Cambridge, MA: MIT Press (2018), 336 pp., $45.00 (cloth). Philosophy of Science, 88(3):562–567, ???? 2022. CODEN PH-SCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Henne:2022:RGD

[123] Céline Henne. Roberto gronda . Dewey’s Philosophy of Science,. Philosophy of Science, 88(3):567–571, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2022:FMb

[124] Anonymous. Front matter. Philosophy of Science, 88(4):??, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Leuschner:2022:HDG

[125] Anna Leuschner and Manuela Fernández Pinto. How dissent on gender bias in academia affects science and society: Learning from the case of climate change denial. Philosophy of Science, 88(4):573–593, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Plaisance:2022:FAB

[126] Kathryn S. Plaisance and Kevin C. Elliott. A framework for analyzing broadly engaged philosophy of science. Philosophy of Science, 88(4):594– 615, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hanley:2022:WCB

[127] Brian J. Hanley. What caused the Bhopal gas tragedy? The philosophical importance of causal and pragmatic details. Philosophy of Science, 88(4): 616–637, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Smeenk:2022:DGR

[128] Chris Smeenk and Christian Wüthrich. Determinism and general relativity. Philosophy of Science, 88(4):638–664, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Arioli:2022:WRQ

[129] Gianni Arioli and Giovanni Valente. What is really quantum in quantum econophysics? Philosophy of Science, 88(4):665–685, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kasirzadeh:2022:NRM

[130] Atoosa Kasirzadeh. A new role for mathematics in empirical sciences. Philosophy of Science, 88(4):686–706, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Babic:2022:ACL

[131] Boris Babic. Approximate coherentism and luck. Philosophy of Science, 88(4):707–725, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fitelson:2022:PCM

[132] Branden Fitelson. A problem for confirmation measure Z. Philosophy of Science, 88(4):726–730, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

McKenna:2022:LMM

[133] Travis McKenna. Lange on minimal model explanations: a defense of Batterman and Rice. Philosophy of Science, 88(4):731–741, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fraser:2022:RMW

[134] Doreen Fraser. Review of Mark Wilson, Physics Avoidance. Philosophy of Science, 88(4):742–750, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Weisberg:2022:EMA

[135] Michael Weisberg. Errol Morris . The Ashtray; or, The Man Who Denied Reality. Philosophy of Science, 88(4):751–754, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dougherty:2022:BRE

[136] John Dougherty. Book review: Elaine Landry, ed. Categories for the Working Philosopher. Philosophy of Science, 88(4):754–757, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Anonymous:2022:FMc

[137] Anonymous. Front matter. Philosophy of Science, 88(5):??, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Potochnik:2022:Pa

[138] Angela Potochnik. Preface. Philosophy of Science, 88(5):759–760, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dinh:2022:CPP

[139] Phuong (Phoebe) Ngoc Dinh and David Danks. Causal pluralism in philosophy: Empirical challenges and alternative proposals. Philosophy of Science, 88(5):761–772, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fischer:2022:CPD

[140] Enno Fischer. Causation and the problem of disagreement. Philosophy of Science, 88(5):773–783, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Casini:2022:VDI

[141] Lorenzo Casini, Alessio Moneta, and Marco Capasso. Variable definition and independent components. Philosophy of Science, 88(5):784–795, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Fang:2022:TMD

[142] Wei Fang. Toward Mechanism 2.1: a dynamic causal approach. Philosophy of Science, 88(5):796–809, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Hunt:2022:UER

[143] Josh Hunt. Understanding and equivalent reformulations. Philosophy of Science, 88(5):810–823, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

DiMarco:2022:WIB

[144] Marina DiMarco. Wishful intelligibility, black boxes, and epidemiological explanation. Philosophy of Science, 88(5):824–834, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Poston:2022:ECI

[145] Ted Poston. Explanatory coherence and the impossibility of confirmation by coherence. Philosophy of Science, 88(5):835–848, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Klein:2022:HAP

[146] Alexander Klein. How American is pragmatism? Philosophy of Science, 88(5):849–859, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Davis:2022:SH

[147] Cruz Austin Davis. Structural humility. Philosophy of Science, 88(5):860– 870, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Brooks:2022:ADC

[148] Daniel S. Brooks. Adaptive design, contingency, and ontological principles for limited beings. Philosophy of Science, 88(5):871–881, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Schneider:2022:CSE

[149] Mike D. Schneider. Creativity in the social epistemology of science. Philosophy of Science, 88(5):882–893, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Reijula:2022:DAT

[150] Samuli Reijula and Jaakko Kuorikoski. The diversity–ability trade-off in scientific problem solving. Philosophy of Science, 88(5):894–905, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Desmond:2022:IRI

[151] Hugh Desmond. Incentivizing replication is insufficient to safeguard default trust. Philosophy of Science, 88(5):906–917, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Miller:2022:WSD

[152] Boaz Miller. When is scientific dissent epistemically inappropriate? Philosophy of Science, 88(5):918–928, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Philippi:2022:MSN

[153] Cristian Larroulet Philippi. On measurement scales: Neither ordinal nor interval? Philosophy of Science, 88(5):929–939, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dresow:2022:MTF

[154] Max Dresow. Measuring time with fossils: a start-up problem in scientific practice. Philosophy of Science, 88(5):940–950, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wu:2022:SDT

[155] Tung-Ying Wu. Structural decision theory. Philosophy of Science, 88(5): 951–960, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pust:2022:DBL

[156] Joel Pust. Dutch books and logical form. Philosophy of Science, 88(5): 961–970, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Meehan:2022:BKP

[157] Alexander Meehan and Snow Zhang. The Borel–Kolmogorov paradox is your paradox too: a puzzle for conditional physical probability. Philosophy of Science, 88(5):971–984, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Haar:2022:SCM

[158] Deborah Haar. Simulation and calibration: Mitigating uncertainty. Philosophy of Science, 88(5):985–996, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dethier:2022:CMI

[159] Corey Dethier. Climate models and the irrelevance of chaos. Philosophy of Science, 88(5):997–1007, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Kawamleh:2022:CML

[160] Suzanne Kawamleh. Can machines learn how clouds work? The epistemic implications of machine learning methods in climate science. Philosophy of Science, 88(5):1008–1020, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Curtis-Trudel:2022:IR

[161] André Curtis-Trudel. Implementation as resemblance. Philosophy of Science, 88(5):1021–1032, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rosenstock:2022:LSD

[162] Sarita Rosenstock. Learning from the shape of data. Philosophy of Science, 88(5):1033–1044, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Rusanen:2022:FDA

[163] Anna-Mari Rusanen, Otto Lappi, Jami Pekkanen, and Jesse Kuokkanen. From fly detectors to action control: Representations in reinforcement learning. Philosophy of Science, 88(5):1045–1054, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Huang:2022:MOS

[164] Linus Ta-Lun Huang, Leonardo Bich, and William Bechtel. Model organisms for studying decision-making: a phylogenetically expanded perspective. Philosophy of Science, 88(5):1055–1066, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Curry:2022:GBM

[165] Devin Sanchez Curry. g as bridge model. Philosophy of Science, 88(5): 1067–1078, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Teira:2022:LCR

[166] David Teira. On the limits of cultural relativism as a debiasing method. Philosophy of Science, 88(5):1079–1089, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Akagi:2022:MOE

[167] Mikio Akagi and Frederick W. Gooding Jr. Microaggressions and objectivity: Experimental measures and lived experience. Philosophy of Science, 88(5):1090–1100, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Pruss:2022:MJD

[168] Dasha Pruss. Mechanical jurisprudence and domain distortion: How predictive algorithms warp the law. Philosophy of Science, 88(5):1101–1112, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wilson:2022:CRP

[169] Alastair Wilson. Counterpossible reasoning in physics. Philosophy of Science, 88(5):1113–1124, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Miller:2022:ICM

[170] Michael E. Miller. Infrared cancellation and measurement. Philosophy of Science, 88(5):1125–1136, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Wells:2022:DCN

[171] Aaron Wells. Du Châtelet on the need for mathematics in physics. Philosophy of Science, 88(5):1137–1148, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Dougherty:2022:SRW

[172] John Dougherty. The substantial role of Weyl symmetry in deriving general relativity from string theory. Philosophy of Science, 88(5):1149–1160, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Lichtenstein:2022:HAH

[173] Eli I. Lichtenstein. How anti-Humeans can embrace a thermodynamic reduction of time’s causal arrow. Philosophy of Science, 88(5):1161–1171, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Chua:2022:NTT

[174] Eugene Y. S. Chua and Craig Callender. No time for time from no-time. Philosophy of Science, 88(5):1172–1184, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Bruner:2022:NBE

[175] Justin P. Bruner. Nash, bargaining and evolution. Philosophy of Science, 88(5):1185–1198, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Aaby:2022:EDN

[176] Bendik Hellem Aaby. The ecological dimension of natural selection. Philosophy of Science, 88(5):1199–1209, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Garson:2022:EGS

[177] Justin Garson. Edmond Goblot’s (1858–1935) selected effects theory of function: a reappraisal. Philosophy of Science, 88(5):1210–1220, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Watkins:2022:EVL

[178] Aja Watkins. The epistemic value of the living fossils concept. Philosophy of Science, 88(5):1221–1233, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

LaCroix:2022:RFR

[179] Travis LaCroix. Reflexivity, functional reference, and modularity: Alternative targets for language origins. Philosophy of Science, 88(5):1234– 1245, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Neal:2022:EOC

[180] Jacob P. Neal. The evolutionary origins of cooperation in the Hominin lineage: a critique of Boyd and Richerson’s cultural group selection account. Philosophy of Science, 88(5):1246–1257, ???? 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

Zerilli:2022:EML

[181] John Zerilli. Explaining machine learning decisions. Philosophy of Science, 89(1):1–19, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/explaining-machine-learning-
decisions/8E20E695A0ADBB2DEC78D0568B78CDF5.

Raz:2022:UDL

[182] Tim Räz. Understanding deep learning with statistical relevance. Philosophy of Science, 89(1):20–41, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/understanding-deep-
learning-with-statistical-relevance/07B47A16041115D1A84C34220B515831.

Grim:2022:STB

[183] Patrick Grim, Frank Seidl, Calum McNamara, Hinton E. Rago, Isabell N. Astor, Caroline Diaso, and Peter Ryner. Scientific theories as Bayesian nets: Structure and evidence sensitivity. Philosophy of Science, 89(1): 42–69, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/scientific-theories-as-bayesian-
nets-structure-and-evidence-sensitivity/9B69390BBAB54D42F6C5CADF93173BFC.

Thagard:2022:ERU

[184] Paul Thagard. Energy requirements undermine substrate independence and mind-body functionalism. Philosophy of Science, 89(1):70– 88, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/energy-requirements-undermine-
substrate-independence-and-mindbody-functionalism/2BB3C2353EFF80F9D5805CDCEA8C3C89.

Colaco:2022:WCM

[185] David Colaço. What counts as a memory? Definitions, hypotheses, and “kinding in progress”. Philosophy of Science, 89(1):89–106, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/what-counts-as-a-memory-definitions-hypotheses-
and-kinding-in-progress/96351DA580FE9D8634AE5E0927E1F819.

Maimon:2022:DNS

[186] Amber Maimon and Meir Hemmo. Does neuroplasticity support the hypothesis of multiple realizability? Philosophy of Science, 89(1): 107–127, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/does-neuroplasticity-
support-the-hypothesis-of-multiple-realizability/6BF636B60D43E8023CD4F24240365B56.

Pettigrew:2022:AFE

[187] Richard Pettigrew. Accuracy-first epistemology without additivity. Philosophy of Science, 89(1):128–151, January 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/accuracyfirst-epistemology-without-additivity/
3EFF180BBFF4CF30426F2DA32E6DB144.

Philippi:2022:TCR

[188] Cristian Larroulet Philippi. There is cause to randomize. Philosophy of Science, 89(1):152–170, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/there-is-cause-to-
randomize/A6238F6192C449095576A6E601296E89.

Dussault:2022:TNE

[189] Antoine C. Dussault. Two notions of ecological function. Philosophy of Science, 89(1):171–179, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/two-notions-
of-ecological-function/0CB394E9D0DD7DD52A0C13D48C63A2E8.

Bauer:2022:AGL

[190] Alexander Max Bauer and Jan Romann. Answers at gunpoint: On Livengood and Sytsma’s revolver case. Philosophy of Science, 89(1):180– 192, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/answers-at-gunpoint-on-livengood-
and-sytsmas-revolver-case/A02B16770F20B1321ACDFEAD44010589. See

[4].

Huss:2022:RWP

[191] John E. Huss. Recent work in the philosophy of medicine: an essay review. Philosophy of Science, 89(1):193–201, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/recent-work-in-the-philosophy-of-medicine-an-
essay-review/72F9BF85D7DA89ED5A52E07DEA6A2EBA.

Feintzeig:2022:BRJ

[192] Benjamin H. Feintzeig. Book review: Jeffrey A. Barrett, The Conceptual Foundations of Quantum Mechanics. Oxford: Oxford University Press (2020), 272 pp., $88.00. Philosophy of Science, 89(1):202–205, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-jeffrey-a-barretts-the-conceptual-
foundations-of-quantum-mechanics-jeffrey-a-barrett-the-conceptual-
foundations-of-quantum-mechanics-oxford-oxford-university-press-
2020-272-pp-8800/5DFC452367BE4EDA57BE0BBC459A0FA1.

Depew:2022:BRJ

[193] David Depew. Book review: Jean Gayon, Interviewed by Victor Petit. Knowledge of Life Today: Conversations on Biology. London & Hoboken, NJ: ISTE/John Wiley and Sons (2019). xx + 378 pp. Philosophy of Science, 89(1):205–208, January 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-jean-gayon-and-victor-petits-knowledge-
of-life-today-jean-gayon-interviewed-by-victor-petit-knowledge-
of-life-today-conversations-on-biology-london-hoboken-nj-istejohn-
wiley-and-sons-2019-xx-378-pp/A1E1175494D1ECC73AA2BCB05589E409.

Anonymous:2022:PVIa

[194] Anonymous. PSA volume 89 issue 1 cover and front matter. Philosophy of Science, 89(1):f1–f4, January 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
1-cover-and-front-matter/FBFF4E92D197659BA5C6309C600DD977.

Anonymous:2022:PVIb

[195] Anonymous. PSA volume 89 issue 1 cover and back matter. Philosophy of Science, 89(2):b1–b2, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
1-cover-and-back-matter/71B010ACC4063F466B059A57299BAC9C.

Fazelpour:2022:DTC

[196] Sina Fazelpour and Daniel Steel. Diversity, trust, and conformity: a simulation study. Philosophy of Science, 89(2):209–231, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/diversity-trust-and-conformity-a-simulation-
study/E1102A15D9D865DD6E988811061EF0CE.

Fuller:2022:EPP

[197] Jonathan Fuller. Epidemics from the population perspective. Philosophy of Science, 89(2):232–251, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/epidemics-from-the-
population-perspective/66249093E4E0502B3E5058AE87FB3249.

Lange:2022:APM

[198] Marc Lange. Against probabilistic measures of explanatory quality. Philosophy of Science, 89(2):252–267, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/against-
probabilistic-measures-of-explanatory-quality/72B1F710A08FF955A0AFEE52889355EB.

Aizawa:2022:AC

[199] Kenneth Aizawa and Drew B. Headley. Abduction and composition. Philosophy of Science, 89(2):268–282, April 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/abduction-and-composition/4ED43ECD77AB946F5542153CC84A4B9F.

Weinberger:2022:SDH

[200] Naftali Weinberger and Colin Allen. Static-dynamic hybridity in dynamical models of cognition. Philosophy of Science, 89(2):283–301, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/staticdynamic-hybridity-in-dynamical-models-
of-cognition/312BBEC4CF032B1ECA7F3E544A5C0384.

Taylor:2022:PEC

[201] Elanor Taylor. Power emergentism and the collapse problem. Philosophy of Science, 89(2):302–318, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/power-emergentism-
and-the-collapse-problem/AEEB06D8D7844DD3B18F52AEAD58D9E6.

Holmes:2022:RCI

[202] Travis Holmes. Reckoning with continuum idealizations: Some lessons from soil hydrology. Philosophy of Science, 89(2):319–336, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/reckoning-with-continuum-idealizations-some-
lessons-from-soil-hydrology/C9D2C2F3186E8A62042342F1F4885D77.

Currie:2022:SMM

[203] Adrian Currie. Speculation made material: Experimental archaeology and maker’s knowledge. Philosophy of Science, 89(2):337–359, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/speculation-made-material-experimental-archaeology-
and-makers-knowledge/17378A91E1D5B829F30BE71FE2A8F578.

Ward:2022:RRP

[204] Barry Ward. Resolving the raven paradox: Simple random sampling, stratified random sampling, and inference to best explanation. Philosophy of Science, 89(2):360–377, April 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/resolving-the-raven-paradox-simple-random-sampling-
stratified-random-sampling-and-inference-to-best-explanation/
FABD2589CAE5EBA50AC826B1F11212D2.

Spanos:2022:STE

[205] Aris Spanos. Severity and trustworthy evidence: Foundational problems versus misuses of frequentist testing. Philosophy of Science, 89(2): 378–397, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/severity-and-trustworthy-evidence-
foundational-problems-versus-misuses-of-frequentist-testing/
91C4FE7B3A14B64720A49EACA4547F28.

Burri:2022:BRR

[206] Susanne Burri. Book review: Richard Pettigrew, Choosing for Changing Selves. Oxford: Oxford University Press (2020), 272 pp., $72.00 (hardcover; also available as an e-book). Philosophy of Science, 89(2): 398–401, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-richard-pettigrews-choosing-
for-changing-selves-richard-pettigrew-choosing-for-changing-
selves-oxford-oxford-university-press-2020-272-pp-7200-hardcover-
also-available-as-an-ebook/AC71919F0D05B7FDB0E037F5F1AAFB2C.

Chalmers:2022:BRG

[207] Alan F. Chalmers. Book review: George E. Smith, and Raghav Seth, Brownian Motion and Molecular Reality: a Study in Theory-Mediated Measurement. Oxford: Oxford University Press (2020), 468 pp., $99.00 (hardcover; also available as an e-book). Philosophy of Science, 89(2):

401–404, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-george-e-smith-and-raghav-
seths-brownian-motion-and-molecular-reality-a-study-in-theorymediated-
measurement-george-e-smith-and-raghav-seth-brownian-motion-and-
molecular-reality-a-study-in-theorymediated-measurement-oxford-
oxford-university-press-2020-468-pp-9900-hardcover-also-available-
as-an-ebook/2289D0DD0C33639DC1D0E6A5188764C6.

Mallon:2022:BRE

[208] Ron Mallon. Book review: Elizabeth Hannon, and Tim Lewens (eds.), Why We Disagree About Human Nature. Oxford: Oxford University Press (2018), 240 pp., $44.95 (hardcover; also available as an e-book). Philosophy of Science, 89(2):405–407, April 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-elizabeth-hannon-and-tim-lewenss-why-
we-disagree-about-human-nature-elizabeth-hannon-and-tim-lewens-
eds-why-we-disagree-about-human-nature-oxford-oxford-university-
press-2018-240-pp-4495-hardcover-also-available-as-an-ebook/
2B2EC3FF7A8C2C35A1702F15E665D737.

Anonymous:2022:RPS

[209] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 89 (2):408–414, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/referees-for-philosophy-of-science/
B5299E8975602CA2A6EE0EC429A306D7.

Anonymous:2022:PVIc

[210] Anonymous. PSA volume 89 issue 2 cover and front matter. Philosophy of Science, 89(2):f1–f4, April 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
2-cover-and-front-matter/CBABBDBE68E875D44674216A376B9426.

Anonymous:2022:PVId

[211] Anonymous. PSA volume 89 issue 2 cover and back matter. Philosophy of Science, 89(3):b1–b2, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
2-cover-and-back-matter/7199E0775ED5473BA56CF7E98E977861.

Sterelny:2022:EAL

[212] Kim Sterelny. Ethnography, archaeology, and the Late Pleistocene. Philosophy of Science, 89(3):415–433, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/ethnography-
archaeology-and-the-late-pleistocene/CDDCDB8DD29F21E5668BC73A0F4E29CE.

Ramsey:2022:EEE

[213] Grant Ramsey and Michael J. Deem. Empathy and the evolutionary emergence of guilt. Philosophy of Science, 89(3):434–453, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/empathy-and-the-evolutionary-emergence-of-guilt/
861EED90948E1575494128853B5B6854.

Hermida:2022:NSI

[214] Margarida Hermida. Natural selection of independently originated life clades. Philosophy of Science, 89(3):454–470, July 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/natural-selection-of-independently-originated-
life-clades/6F479CCA0FF6089AF947F2AC04C8CB65.

Guala:2022:ROI

[215] Francesco Guala. Rescuing ontological individualism. Philosophy of Science, 89(3):471–485, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/rescuing-ontological-
individualism/97C99E21FB5B664591FD86ACCDFA5D49.

Calosi:2022:BSA

[216] Claudio Calosi. The bound state answer to the special composition question. Philosophy of Science, 89(3):486–503, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/bound-state-answer-to-the-special-composition-
question/FC664075A884F8231C788BCD829F7C1E.

Stemeroff:2022:SPM

[217] Noah Stemeroff. Scientific perspectivism and the methodology of modern mathematical physics. Philosophy of Science, 89(3):504–520, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/scientific-perspectivism-and-the-methodology-
of-modern-mathematical-physics/1EABE7AC1DB8805B9DD14140247FA46A.

u:2022:PI

[218] Ana-Maria Cre u. Perspectival instruments. Philosophy of Science, 89(3): 521–541, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/perspectival-instruments/611F11C0C8BC865FA53B732FA92481A3.

Schindler:2022:TVD

[219] Samuel Schindler. Theoretical virtues: Do scientists think what philosophers think they ought to think? Philosophy of Science, 89(3):542–564, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/theoretical-
virtues-do-scientists-think-what-philosophers-think-they-ought-
to-think/0C25D6844ADF529BB2912DD5DEF901CE.

Dupre:2022:ROV

[220] Gabe Dupre. Realism and observation: The view from generative grammar. Philosophy of Science, 89(3):565–584, July 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/realism-and-observation-the-view-from-generative-
grammar/66344E922DFAB0F861842509F143EA03.

Dewulf:2022:BHR

[221] Fons Dewulf. Beyond Hempel: Reframing the debate about scientific explanation. Philosophy of Science, 89(3):585–603, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/beyond-hempel-reframing-the-debate-about-scientific-
explanation/25F4BE0A06F0A4370615131ECEB156C6.

Stewart:2022:DSP

[222] Rush T. Stewart and Michael Nielsen. Distention for sets of probabilities. Philosophy of Science, 89(3):604–620, July 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/distention-for-sets-of-probabilities/C67F8F907D36C975BEB0B6695DFFDA74.

Nielsen:2022:BAA

[223] Michael Nielsen. On the best accuracy arguments for probabilism. Philosophy of Science, 89(3):621–630, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/on-the-best-
accuracy-arguments-for-probabilism/EC933C585F4849477A42114040A5C849.

Wagner:2022:PRJ

[224] Carl G. Wagner. Postscript to Richard Jeffrey’s Conditioning, Kinematics, and Exchangeability. Philosophy of Science, 89(3):631–643, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/postscript-to-richard-jeffreys-conditioning-
kinematics-and-exchangeability/37C22FA36A48C0E54432FD11F6D99497.

Barrett:2022:CSC

[225] Thomas William Barrett. Coordinates, structure, and classical mechanics: a review of Jill North’s Physics, Structure, and Reality. Philosophy of Science, 89(3):644–653, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/coordinates-
structure-and-classical-mechanics-a-review-of-jill-norths-physics-
structure-and-reality/FCA25B3ECD7892E07F7912DD5072CDC5.

Crasnow:2022:BRM

[226] Sharon Crasnow. Book review: Maya J. Goldenberg, Vaccine Hesitancy: Public Trust, Expertise, and the War on Science. Pittsburgh: University of Pittsburgh Press (2021), 264 pp., $32.00 (paperback). Philosophy of Science, 89(3):654–658, July 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-maya-j-goldenbergs-vaccine-hesitancy-
public-trust-expertise-and-the-war-on-science-maya-j-goldenberg-
vaccine-hesitancy-public-trust-expertise-and-the-war-on-science-
pittsburgh-university-of-pittsburgh-press-2021-264-pp-3200-paperback/
6EECF293308C213C7FD47CEBD5166C42.

Veit:2022:BRP

[227] Walter Veit. Book review: Peter Godfrey-Smith, Metazoa: Animal Minds and the Birth of Consciousness. Glasgow: William Collins (2020), 288 pp., $24.99 (hardcover; also available in paperback, nook, and audiobook formats). Philosophy of Science, 89(3):658– 660, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-peter-godfreysmiths-
metazoa-animal-minds-and-the-birth-of-consciousness-peter-godfreysmith-
metazoa-animal-minds-and-the-birth-of-consciousness-glasgow-
william-collins-2020-288-pp-2499-hardcover-also-available-in-
paperback-nook-and-audiobook-formats/E1B157CECB3A5EDF21A63165067A7C05.

Anonymous:2022:PVIe

[228] Anonymous. PSA volume 89 issue 3 cover and front matter. Philosophy of Science, 89(3):f1–f4, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
3-cover-and-front-matter/54FE624AE1A1D1F330FA7054DD866CE6.

Anonymous:2022:PVIf

[229] Anonymous. PSA volume 89 issue 3 cover and back matter. Philosophy of Science, 89(3):b1–b2, July 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
3-cover-and-back-matter/913DFDF2A01664C3256B51BFD8F33B25.

Fagerberg:2022:WMD

[230] Harriet Fagerberg. Why mental disorders are not like software bugs. Philosophy of Science, 89(4):661–682, October 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/why-mental-disorders-are-not-like-software-bugs/
465B1BE67DE7749A9E9A54F67506A5A7.

Jonkisz:2022:WMB

[231] Jakub Jonkisz. What makes behavioral measures of consciousness subjective and direct? Philosophy of Science, 89(4):683–700, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/what-makes-behavioral-measures-of-consciousness-
subjective-and-direct/199E2647B132B907AF8054816BF3BD22.

Hacohen:2022:WNR

[232] Ori Hacohen. What are neural representations? A Cummins functions approach. Philosophy of Science, 89(4):701–720, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/what-are-neural-representations-a-cummins-functions-
approach/B82AA1FC816ED3CBD53DC3B631390B1E.

Nersessian:2022:REP

[233] Nancy J. Nersessian and Miles MacLeod. Rethinking ethnography for philosophy of science. Philosophy of Science, 89(4):721–741, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/rethinking-ethnography-for-philosophy-of-science/
98185690361DB609B26B95C11C2504E7.

McCullough-Benner:2022:MRM

[234] Colin McCullough-Benner. The metarepresentational role of mathematics in scientific explanations. Philosophy of Science, 89(4):742–760, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/metarepresentational-role-of-mathematics-in-
scientific-explanations/F2CA15A4C503EB51EFB86961ADE08E6E.

Ghomi:2022:SDL

[235] Hamed Tabatabaei Ghomi. Setting the demons loose: Computational irreducibility does not guarantee unpredictability or emergence. Philosophy of Science, 89(4):761–783, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/setting-the-
demons-loose-computational-irreducibility-does-not-guarantee-
unpredictability-or-emergence/89063B3CC696E6CF03ECD99888BE7124.

Ehrlich:2022:PNU

[236] Philip Ehrlich. Are points (necessarily) unextended? Philosophy of Science, 89(4):784–801, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/are-points-necessarily-
unextended/58B5F3EE299B4808B7BCA5A858A0C81A.

Lloyd:2022:VDC

[237] Elisabeth Lloyd, Greg Lusk, Stuart Gluck, and Seth McGinnis. Varieties of data-centric science: Regional climate modeling and model organism research. Philosophy of Science, 89(4):802–823, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/varieties-of-datacentric-science-regional-climate-
modeling-and-model-organism-research/FFD43887EFECFE33AEADDDFA7FD35048.

DiBella:2022:FIL

[238] Nicholas DiBella. Fair infinite lotteries, qualitative probability, and regularity. Philosophy of Science, 89(4):824–844, October 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/fair-infinite-lotteries-qualitative-probability-
and-regularity/88F15215C3FB30BA3C642C95DFB87797.

Kinney:2022:WAW

[239] David Kinney. Why average when you can stack? Better methods for generating accurate group credences. Philosophy of Science, 89(4): 845–863, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/why-average-when-you-
can-stack-better-methods-for-generating-accurate-group-credences/
7E57997732ABB091621D010AC7DE3CF2.

Hubert:2022:RAN

[240] Mario Hubert. Review of Alyssa Ney’s The World in the Wave Function: a Metaphysics for Quantum Physics. New York: Oxford University Press (2021), 288 pp., $35.00 (hardcover). Philosophy of Science, 89(4): 864–868, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-alyssa-neys-the-world-
in-the-wave-function-a-metaphysics-for-quantum-physics-alyssa-
ney-the-world-in-the-wave-function-a-metaphysics-for-quantum-
physics-new-york-oxford-university-press-2021-288-pp-3500-hardcover/
F54D1DE0E70220C8516144A3B5961EBC.

Wilson:2022:RAC

[241] Joseph Wilson. Review of Adrian Currie’s Rock, Bone, and Ruin: an Optimist’s Guide to the Historical Sciences. Cambridge, MA: MIT Press (2018), 372 pp., $35.00 (hardcover). Philosophy of Science, 89(4):868– 871, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-adrian-curries-rock-
bone-and-ruin-an-optimists-guide-to-the-historical-sciences-
adrian-currie-rock-bone-and-ruin-an-optimists-guide-to-the-historical-
sciences-cambridge-ma-mit-press-2018-372-pp-3500-hardcover/CC9C0DCB32E1D0B0BB207BBAF86E5215.

Parke:2022:RCC

[242] Emily C. Parke. Review of Carol E. Cleland’s The Quest for a Universal Theory of Life: Searching for Life as We Don’t Know It Cambridge: Cambridge University Press (2019), 258 pp., $79.99 (hardcover). Philosophy of Science, 89(4):872–875, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/review-of-carol-
e-clelands-the-quest-for-a-universal-theory-of-life-searching-
for-life-as-we-dont-know-it-carol-e-cleland-the-quest-for-a-
universal-theory-of-life-searching-for-life-as-we-dont-know-
it-cambridge-cambridge-university-press-2019-258-pp-7999-hardcover/
3E2C0EB4BED6320C4A1AC359224356F4.

Anonymous:2022:PVIg

[243] Anonymous. PSA volume 89 issue 4 cover and front matter. Philosophy of Science, 89(4):f1–f4, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
4-cover-and-front-matter/F1B7EA86059E0642D773ED6F87953F7C.

Anonymous:2022:PVIh

[244] Anonymous. PSA volume 89 issue 4 cover and back matter. Philosophy of Science, 89(4):b1–b2, October 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-89-issue-
4-cover-and-back-matter/32D87E08DCC0F2E41D2F440CB46C7E90.

Potochnik:2022:Pb

[245] Angela Potochnik. Preface. Philosophy of Science, 89(5):877–878, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/preface/D1D52F69AB867110C6E4A7F4423348F9.

Rowbottom:2022:CMS

[246] Darrell P. Rowbottom. Can meaningless statements be approximately true? On relaxing the semantic component of scientific realism. Philosophy of Science, 89(5):879–888, December 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/can-meaningless-statements-be-approximately-
true-on-relaxing-the-semantic-component-of-scientific-realism/
3D110B7E9723D7AC0DE4B69B19FDE71A.

Ross:2022:IOH

[247] Lauren N. Ross and James F. Woodward. Irreversible (one-hit) and reversible (sustaining) causation. Philosophy of Science, 89(5):889–898, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/irreversible-onehit-and-reversible-
sustaining-causation/13A291081C0E09D3C435436916C7D0AD.

Sober:2022:PSS

[248] Elliott Sober. Philosophy in science: Some personal reflections. Philosophy of Science, 89(5):899–907, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/philosophy-in-science-
some-personal-reflections/BC6A652737B61B992038A4135628ECD8.

Knuuttila:2022:EPS

[249] Tarja Knuuttila and Andrea Loettgers. (Un)Easily possible synthetic biology. Philosophy of Science, 89(5):908–917, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/uneasily-possible-synthetic-biology/4A099D89C379CEA01BC05176499CEBBE.

Verreault-Julien:2022:RNA

[250] Philippe Verreault-Julien. Representing non-actual targets? Philosophy of Science, 89(5):918–927, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/representing-nonactual-
targets/6B88D14BB2B8ADB310B7490ED6D54CD9.

Bartha:2022:PAC

[251] Paul Bartha. Prospects for analogue confirmation. Philosophy of Science, 89(5):928–938, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/prospects-for-analogue-
confirmation/B3582831F209E716C5A6D7630F9643DD.

Steele:2022:TEA

[252] Katie Steele and H. Orri Stefánsson. Transformative experience, awareness growth, and the limits of rational planning. Philosophy of Science, 89(5): 939–948, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/transformative-experience-
awareness-growth-and-the-limits-of-rational-planning/351EA830C0CAA699A24F58408B66FABF.

Ghomi:2022:CCO

[253] Hamed Tabatabaei Ghomi and Jacob Stegenga. Conventional choices in outcome measures influence meta-analytic results. Philosophy of Science, 89(5):949–959, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/conventional-choices-
in-outcome-measures-influence-metaanalytic-results/A967D5FC97C44E429899417E602D4024.

Fletcher:2022:RMA

[254] Samuel C. Fletcher. Replication is for meta-analysis. Philosophy of Science, 89(5):960–969, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/replication-
is-for-metaanalysis/AE2F21B8AEF1ABB64C9AE26EA827C382.

Holman:2022:WFC

[255] Bennett Holman. What FiveThirtyEight can teach us about solving the replication crisis. Philosophy of Science, 89(5):970–979, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/what-fivethirtyeight-can-teach-us-about-solving-
the-replication-crisis/237AF4E17114FC72ADA53334ABD17543.

Kovaka:2022:MAC

[256] Karen Kovaka. Meta-analysis and conservation science. Philosophy of Science, 89(5):980–990, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/metaanalysis-and-
conservation-science/971DEE9922BF268456150D56DCCDDB2C.

Leonelli:2022:OSE

[257] Sabina Leonelli. Open science and epistemic diversity: Friends or foes? Philosophy of Science, 89(5):991–1001, December 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/open-science-and-epistemic-diversity-friends-
or-foes/6206992DB99036C83549E1BE5348E04E.

Lee:2022:CAE

[258] Carole J. Lee. Certified amplification: an emerging scientific norm and ethos. Philosophy of Science, 89(5):1002–1012, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/certified-amplification-an-emerging-scientific-
norm-and-ethos/C463A7C3DB30748CAD4DB492709A3324.

Elliott:2022:OSN

[259] Kevin C. Elliott. Open science for non-specialists: Making open science meaningful beyond the scientific community. Philosophy of Science, 89(5): 1013–1023, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/open-science-for-nonspecialists-
making-open-science-meaningful-beyond-the-scientific-community/
70855E04352B57CA03319FE385B6D095.

Bobadilla:2022:WSF

[260] Hernán Bobadilla. What is at stake in the formalization of a chronostratigraphic unit? A case study on the Anthropocene. Philosophy of Science, 89(5):1024–1033, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/what-is-at-stake-
in-the-formalization-of-a-chronostratigraphic-unit-a-case-study-
on-the-anthropocene/289C9814E553A3DB39AAF33261CFD737.

Schroeder:2022:LDS

[261] S. Andrew Schroeder. The limits of democratizing science: When scientists should ignore the public. Philosophy of Science, 89(5):1034–1043, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/limits-of-democratizing-science-
when-scientists-should-ignore-the-public/0722C3510F730BA972AAC17FE53CFFE8.

Slater:2022:TSP

[262] Matthew H. Slater and Emily R. Scholfield. Trust of science as a public collective good. Philosophy of Science, 89(5):1044–1053, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/trust-of-science-as-a-public-collective-good/
A3B36A4FFE2535A0218D6B61BD6D5B5A.

Nyrup:2022:LVT

[263] Rune Nyrup. The limits of value transparency in machine learning. Philosophy of Science, 89(5):1054–1064, December 2022. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/limits-of-value-transparency-in-machine-learning/
9221D6E444184FD989E17651DBA29F2C.

Sullivan:2022:IRU

[264] Emily Sullivan. Inductive risk, understanding, and opaque machine learning models. Philosophy of Science, 89(5):1065–1074, December 2022. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/inductive-risk-understanding-and-opaque-machine-
learning-models/DB9D3522CEA45E62C98BA640DAAAC95E.

Fleisher:2023:RCE

[265] Will Fleisher and Dunja Seselja. Responsibility for collective epistemic harms. Philosophy of Science, 90(1):1–20, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/responsibility-for-collective-epistemic-harms/
A92EE23F12E63D85381CDD3C36F4DB52.

Neto:2023:CME

[266] Celso Neto and W. Ford Doolittle. A chemostat model for evolution by persistence: Clade selection and its explanatory autonomy. Philosophy of Science, 90(1):21–38, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/chemostat-model-
for-evolution-by-persistence-clade-selection-and-its-explanatory-
autonomy/976D7302BADAD3339B2A9D3A01738AD2.

Krickel:2023:UMW

[267] Beate Krickel. The unconscious mind worry : a mechanistic-explanatory strategy. Philosophy of Science, 90(1):39–59, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/unconscious-mind-worry-a-mechanisticexplanatory-
strategy/F438C35F601657F5D42FCD6C583DC35E.

Hoek:2023:FCO

[268] Daniel Hoek. Forced changes only: a new take on the law of inertia. Philosophy of Science, 90(1):60–76, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/forced-changes-only-a-new-take-on-the-law-of-
inertia/3B1DE948FEEC58CD6EB3627F7976CB8F.

Stradis:2023:BCH

[269] Athamos Stradis. Backward causation: Harder than it looks. Philosophy of Science, 90(1):77–91, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/backward-causation-
harder-than-it-looks/EF80FDC7A690345E6CA5A58D05AF2298.

Kelley:2023:ACI

[270] Mikayla Kelley. On accuracy and coherence with infinite opinion sets. Philosophy of Science, 90(1):92–128, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/on-accuracy-and-coherence-with-infinite-opinion-
sets/6B348D5FC458630398D829263198F3BF.

Roche:2023:PPM

[271] William Roche and Elliott Sober. Purely probabilistic measures of explanatory power: a critique. Philosophy of Science, 90(1):129–149, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/purely-probabilistic-measures-
of-explanatory-power-a-critique/57A2249480FB318E62492C357B9E91A0.

Friend:2023:HHA

[272] Toby Friend. How to be Humean about idealization laws. Philosophy of Science, 90(1):150–170, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/how-to-be-humean-
about-idealization-laws/46DE07C4EC8726A857C574C8196CCEB7.

Duguid:2023:SHM

[273] Callum Duguid. Symmetries as Humean metalaws. Philosophy of Science, 90(1):171–187, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/symmetries-as-humean-
metalaws/6C71BBA28635D16DAF80BF95032877BE.

Dewar:2023:ISC

[274] Neil Dewar. On internal structure, categorical structure, and representation. Philosophy of Science, 90(1):188–195, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/on-internal-structure-categorical-structure-
and-representation/AF4CA5E4A8C4E57D9C8CD8753F69DA71.

Wray:2023:TKH

[275] K. Brad Wray. Thomas Kuhn, hyperbole, and the ashtray: Evidence of Morris’ faulty memory. Philosophy of Science, 90(1):196– 199, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/thomas-kuhn-hyperbole-and-the-
ashtray-evidence-of-morris-faulty-memory/C9E588E47B35E4F938FF4CE59EA9AAF1.

Paslaru:2023:NTT

[276] Viorel Pâslaru. New textbooks for teaching philosophy of science. Philosophy of Science, 90(1):200–208, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/new-textbooks-for-teaching-philosophy-of-science/
A43AA3F69AA7CE22FE4F0DBAD00A903D.

Buskell:2023:HTG

[277] Andrew Buskell, Edwin Etieyibo, Catherine Kendig, Raphael Uchôa, and Robert A. Wilson. How to think with the global south. Philosophy of Science, 90(1):209–217, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/how-to-think-with-
the-global-south/9AEE8FBF2EED925B01730F59F8805F18.

Virmajoki:2023:RPR

[278] Veli Virmajoki. Review of Paul A. Roth’s The Philosophical Structure of Historical Explanation — Paul A. Roth, The Philosophical Structure of Historical Explanation . Evanston, IL: Northwestern University Press (2020), 208 pp. $34.95 (paperback). Philosophy of Science, 90(1):218– 221, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-paul-a-roths-the-philosophical-
structure-of-historical-explanation-paul-a-roth-the-philosophical-
structure-of-historical-explanation-evanston-il-northwestern-
university-press-2020-208-pp-3495-paperback/3F91E3462834E56669B4379A11A819D2.

Ben-Menahem:2023:RWD

[279] Yemima Ben-Menahem. Review of William Demopoulos’ On Theories: Logical Empiricism and the Methodology of Modern Physics — William Demopoulos, On Theories: Logical Empiricism and the Methodology of Modern Physics , edited by Michael Friedman. Cambridge: Harvard University Press (2022), 272 pp., $39.95 (hardcover). Philosophy of Science, 90(1):222–225, January 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-william-demopoulos-on-theories-logical-
empiricism-and-the-methodology-of-modern-physics-william-demopoulos-
on-theories-logical-empiricism-and-the-methodology-of-modern-
physics-edited-by-michael-friedman-cambridge-harvard-university-
press-2022-272-pp-3995-hardcover/4E3E5F4BC3B747E4E21BE521E3E4BCB8.

Anonymous:2023:PVIa

[280] Anonymous. PSA volume 90 issue 1 cover and front matter. Philosophy of Science, 90(1):f1–f4, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
1-cover-and-front-matter/E0244F2A8C4FFE24A3BDD7C05A8F9CD6.

Anonymous:2023:PVIb

[281] Anonymous. PSA volume 90 issue 1 cover and back matter. Philosophy of Science, 90(1):b1–b2, January 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
1-cover-and-back-matter/34E2C92FFBB1E41EAD19473C5952CFDF.

Zhao:2023:MNV

[282] Kino Zhao. Measuring the nonexistent: Validity before measurement. Philosophy of Science, 90(2):227–244, April 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/measuring-the-nonexistent-validity-before-measurement/
5F99E48B565964016D15956443E01316.

Kostic:2023:DTE

[283] Daniel Kostić and Kareem Khalifa. Decoupling topological explanations from mechanisms. Philosophy of Science, 90(2):245–268, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/decoupling-topological-explanations-from-mechanisms/
4801B5E2E1B920E818D5F27CEABA84BE.

Seifert:2023:CBR

[284] Vanessa A. Seifert. The chemical bond is a real pattern. Philosophy of Science, 90(2):269–287, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/chemical-bond-is-
a-real-pattern/32016AD6358F5C350336064877B45FAF.

Neth:2023:DSP

[285] Sven Neth. A dilemma for Solomonoff prediction. Philosophy of Science, 90(2):288–306, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/dilemma-for-
solomonoff-prediction/237324EF10CB4EA83841185DD1EFE50C.

Freeborn:2023:INS

[286] David Peter Wallis Freeborn. The invention of new strategies in bargaining games. Philosophy of Science, 90(2):307–335, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/invention-of-new-strategies-in-bargaining-games/
1A1D148E1557E606F404BAB6707B361B.

Ventura:2023:SIC

[287] Rafael Ventura. Structural inequality in collaboration networks. Philosophy of Science, 90(2):336–353, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/structural-inequality-
in-collaboration-networks/8EED7D94EAC5448F910323A2DE7839AC.

Haueis:2023:ECF

[288] Philipp Haueis. Exploratory concept formation and tool development in neuroscience. Philosophy of Science, 90(2):354–375, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/exploratory-concept-formation-and-tool-development-
in-neuroscience/4AA9B97C960E54D95A9F8E68B27FB47D.

Imbert:2023:FVS

[289] Cyrille Imbert and Vincent Ardourel. Formal verification, scientific code, and the epistemological heterogeneity of computational science. Philosophy of Science, 90(2):376–394, April 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/formal-verification-scientific-code-and-the-
epistemological-heterogeneity-of-computational-science/60D6DDBEB759DC87EB49AB521BA1F222.

Dong:2023:WDI

[290] Zili Dong. Well-defined interventions and causal variable choice. Philosophy of Science, 90(2):395–412, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/welldefined-
interventions-and-causal-variable-choice/F6C8401DDDEF11816B04D69FA19F4F0D.

Koskinen:2023:PO

[291] Inkeri Koskinen. Participation and objectivity. Philosophy of Science, 90 (2):413–432, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/participation-and-objectivity/
B4B7A77B16B40C64CEFB5336B81881B8.

Morrow:2023:CRA

[292] Katie H. Morrow. A causal-role account of ecological role functions. Philosophy of Science, 90(2):433–453, April 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/causalrole-account-of-ecological-role-functions/
AB789914A413A25C0917BFE606FE520C.

Koberinski:2023:LEF

[293] Adam Koberinski and Chris Smeenk. Λ and the limits of effective field theory. Philosophy of Science, 90(2):454–474, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/and-the-limits-of-effective-field-theory/B420BA49A7548036DBDA3D0A13119B29.

Andersen:2023:RJW

[294] H. K. Andersen. Review of James Woodward’s Causation with a Human Face — James Woodward, Causation with a Human Face. New York, NY: Oxford University Press (2021), 424 pp., $99.00 (hardcover). Philosophy of Science, 90(2):475–479, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/review-of-james-
woodwards-causation-with-a-human-face-james-woodward-causation-
with-a-human-face-new-york-ny-oxford-university-press-2021-424-
pp-9900-hardcover/00DD3AB08129BDF6A364A676C2ADB0D9.

Franco:2023:RMJ

[295] Paul L. Franco. Review of Matthew J. Brown’s Science and the Moral Imagination: a New Ideal for Values in Science — Matthew J. Brown, Science and Moral Imagination: a New Ideal for Values in Science. With a foreword by Kim Stanley Robinson. Pittsburgh: Pittsburgh University Press (2020), 288 pp., $50.00 (Hardcover). Available open access with supplementary materials: https://valuesinscience.com. Philosophy of Science, 90(2):479–483, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-matthew-j-browns-science-and-the-moral-
imagination-a-new-ideal-for-values-in-science-matthew-j-brown-
science-and-moral-imagination-a-new-ideal-for-values-in-science-
with-a-foreword-by-kim-stanley-robinson-pittsburgh-pittsburgh-
university-press-2020-288-pp-5000-hardcover-available-open-access-
with-supplementary-materials-httpsvaluesinsciencecom/6C68099C80C137286EB92955CED65E3F.

Anonymous:2023:RPS

[296] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 90 (2):484–488, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/referees-for-philosophy-of-science/
9167C5ECFB5C737E4D28C1B5E945D162.

Anonymous:2023:PVIc

[297] Anonymous. PSA volume 90 issue 2 cover and front matter. Philosophy of Science, 90(2):f1–f4, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
2-cover-and-front-matter/C587AB3D19D9292D74CF7B6B66F394B7.

Anonymous:2023:PVId

[298] Anonymous. PSA volume 90 issue 2 cover and back matter. Philosophy of Science, 90(2):b1–b2, April 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
2-cover-and-back-matter/4300D8D22C95916C5970EFD4B83C2137.

Bauer:2023:PBS

[299] Alexander Max Bauer and Stephan Kornmesser. Poisoned babies, shot fathers, and ruined experiments: Experimental evidence in favor of the compositionality constraint of actual causation. Philosophy of Science, 90(3):489–517, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/poisoned-babies-shot-
fathers-and-ruined-experiments-experimental-evidence-in-favor-
of-the-compositionality-constraint-of-actual-causation/4D0C683AF6A00F0BA2B2A0923256166F.

Stuart:2023:SEC

[300] Michael T. Stuart. Scientists are epistemic consequentialists about imagination. Philosophy of Science, 90(3):518–538, July 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/scientists-are-epistemic-consequentialists-about-
imagination/90B0D0CEDABE9104F65A57B15E547F4F.

Schiller:2023:GII

[301] Henry Ian Schiller. Genericity and inductive inference. Philosophy of Science, 90(3):539–557, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/genericity-and-inductive-
inference/247948E5B6625D68EEC45554E84FD550.

Wills:2023:RM

[302] James Wills. Representing mixtures. Philosophy of Science, 90(3): 558–579, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/representing-mixtures/63E94B24A8F3E01838E43615B8F99F5A.

Bar-On:2023:CRC

[303] Kati Kish Bar-On. Connecting the revolutionary with the conventional: Rethinking the differences between the works of Brouwer, Heyting, and Weyl. Philosophy of Science, 90(3):580–602, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/connecting-the-revolutionary-with-the-conventional-
rethinking-the-differences-between-the-works-of-brouwer-heyting-
and-weyl/9376F606893AEDB9AD887548C59D6099.

Sakamoto:2023:LBW

[304] Takahiro Sakamoto. Is lottery a better way of resource distribution than baseline funding? Philosophy of Science, 90(3):603–628, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/is-lottery-a-better-way-of-resource-distribution-
than-baseline-funding/183F013DFD2FA45F7CA7A542DB6660A2.

Antoniou:2023:RDM

[305] Antonis Antoniou. Robustness and dark-matter observation. Philosophy of Science, 90(3):629–647, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/robustness-and-darkmatter-
observation/D53AD4A11B16B999B03FC8A5129288BE.

Fatollahi:2023:SSF

[306] Alireza Fatollahi and Kasra Alishahi. Simplicity and the sub-family problem for model selection. Philosophy of Science, 90(3):648–668, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/simplicity-and-the-subfamily-problem-for-model-
selection/E7899A5499C162E1B281EB9E181B0CF7.

Murphy:2023:FCD

[307] Alice Murphy. Form and content: a defense of aesthetic value in science. Philosophy of Science, 90(3):669–685, July 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/form-and-content-a-defense-of-aesthetic-value-
in-science/469FEB52B2C8E63BE25CE562C942C2C8.

Conley:2023:DD

[308] Brandon A. Conley. Dispositionalism and dysfunction. Philosophy of Science, 90(3):686–703, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/dispositionalism-
and-dysfunction/1F7F03896C61CE9AFDE93D916561E45E.

Bourrat:2023:PFG

[309] Pierrick Bourrat. A Pricean formalization of Gaia. Philosophy of Science, 90(3):704–720, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/pricean-formalization-
of-gaia/8458502A3973CDCED5EF353AE8895D67.

Thompson:2023:CMR

[310] Morgan Thompson. Characterizing and measuring racial discrimination in public health research. Philosophy of Science, 90(3):721–743, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/characterizing-and-measuring-racial-discrimination-
in-public-health-research/7D7EB934D1E83F71A92612D0DAE1AF77.

Dethier:2023:SSM

[311] Corey Dethier. Supposition and (statistical) models. Philosophy of Science, 90(3):744–749, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/supposition-and-
statistical-models/5BADC6B3078112FE28B94DD1BA005EA9.

Hazelwood:2023:RCH

[312] Caleb Hazelwood. Review of Charles H. Pence’s The Causal Structure of Natural Selection — Charles H. Pence, The Causal Structure of Natural Selection. Elements in the Philosophy of Biology. Cambridge: Cambridge University Press (2021), 75 pp. $22.00 (paperback). Philosophy of Science, 90(3):750–753, July 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-charles-h-pences-the-causal-structure-
of-natural-selection-charles-h-pence-the-causal-structure-of-
natural-selection-elements-in-the-philosophy-of-biology-cambridge-
cambridge-university-press-2021-75-pp-2200-paperback/E86C902DF8319C0D46FE5F3A585B228E.

Favela:2023:RGP

[313] Luis H. Favela. Review of Gualtiero Piccinini’s Neurocognitive Mechanisms: Explaining Biological Cognition — Gualtiero Piccinini, Neurocognitive Mechanisms: Explaining Biological Cognition. Oxford: Oxford University Press (2020), 416 pp. $115.00 (hardcover). Philosophy of Science, 90(3):753–756, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/review-of-gualtiero-piccininis-
neurocognitive-mechanisms-explaining-biological-cognition-gualtiero-
piccinini-neurocognitive-mechanisms-explaining-biological-cognition-
oxford-oxford-university-press-2020-416-pp-11500-hardcover/A4FAB2C886439A395CA248CF39F36155.

Hicks:2023:RNC

[314] Daniel J. Hicks. Review of Nancy Cartwright’s A Philosopher Looks at Science — Nancy Cartwright, A Philosopher Looks at Science. Cambridge: Cambridge University Press (2022), 210 pp. $12.95 (paperback). Philosophy of Science, 90(3):756–760, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/review-of-nancy-

cartwrights-a-philosopher-looks-at-science-nancy-cartwright-
a-philosopher-looks-at-science-cambridge-cambridge-university-
press-2022-210-pp-1295-paperback/1BFFFC2789221D7FF1F03C6790825E97.

Anonymous:2023:PVIe

[315] Anonymous. PSA volume 90 issue 3 cover and front matter. Philosophy of Science, 90(3):f1–f4, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
3-cover-and-front-matter/A03C9D8ABB80C0B60A020DC03E1296C3.

Anonymous:2023:PVIf

[316] Anonymous. PSA volume 90 issue 3 cover and back matter. Philosophy of Science, 90(3):b1–b2, July 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
3-cover-and-back-matter/15826EC9117E47955B052E93D600D718.

Dyck:2023:NMB

[317] Keith Dyck. On no-miracles and the base-rate fallacy. Philosophy of Science, 90(4):761–776, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/on-nomiracles-and-
the-baserate-fallacy/789FFB92DAAF1A00E58F1FDFF86B34E5.

Karaca:2023:NAM

[318] Koray Karaca. A network account of models in high energy physics experiments. Philosophy of Science, 90(4):777–796, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/network-account-of-models-in-high-energy-physics-
experiments/B12998DA7E3ED51254835417DE048C67.

Jacobs:2023:NCN

[319] Caspar Jacobs. The nature of a constant of nature: The case of G. Philosophy of Science, 90(4):797–816, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/nature-of-a-constant-of-nature-the-case-of-g/
4FF66557DB8796B59826E6B2C088C9E2.

Fischborn:2023:ERM

[320] Marcelo Fischborn. The evolutionary roots of moral responsibility. Philosophy of Science, 90(4):817–835, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/evolutionary-roots-of-moral-responsibility/9304FAE36ED9FB8D55BD03002472903D.

Jhun:2023:MMR

[321] Jennifer S. Jhun. Multi-model reasoning in economics: The case of COMPASS. Philosophy of Science, 90(4):836–854, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/multimodel-reasoning-in-economics-the-case-of-
compass/3502EB5E7CD5CF6CF4AC4ED9F3C2DA8C.

Adlam:2023:TAI

[322] Emily Adlam. The temporal asymmetry of influence is not statistical. Philosophy of Science, 90(4):855–879, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/temporal-asymmetry-of-influence-is-not-statistical/
E891E4BB996C75BCC819EAF2A82648FE.

Nefdt:2023:MSM

[323] Ryan M. Nefdt. Motivating a scientific modeling continuum: The case of “Natural models” in the COVID-19 pandemic. Philosophy of Science, 90 (4):880–900, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/motivating-a-scientific-
modeling-continuum-the-case-of-natural-models-in-the-covid19-
pandemic/C02E00D64CACD643418E2A1B8320AF76.

Khalili:2023:DAR

[324] Mahdi Khalili. A dialogue among recent views of entity realism. Philosophy of Science, 90(4):901–921, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/dialogue-among-recent-views-of-entity-realism/
1AD9551D0CB7D8665EAC1E937F5C03DA.

Lorenzetti:2023:FRL

[325] Lorenzo Lorenzetti. Functionalism, reductionism, and levels of reality. Philosophy of Science, 90(4):922–936, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/functionalism-reductionism-and-levels-of-reality/
C5BF00ECAF75F2B9C54EB3E25EA6CA63.

Friedman:2023:SDF

[326] Daniel C. Friedman and Dunja Seselja. Scientific disagreements, fast science and higher-order evidence. Philosophy of Science, 90(4):937– 957, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/scientific-disagreements-fast-
science-and-higherorder-evidence/59AD3DE35564D9E3774D11C1FD051A6D.

Brenner:2023:CPS

[327] Andrew Brenner. Conditional probabilities and symmetric grounding. Philosophy of Science, 90(4):958–973, October 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/conditional-probabilities-and-symmetric-grounding/
1D443BA41D32AE9A0F33A50D5E6246F3.

Schmitz:2023:EUP

[328] Timothy Schmitz. On epistemically useful physical computation. Philosophy of Science, 90(4):974–984, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/on-
epistemically-useful-physical-computation/1EB8EA626478CFDD6636A06F73662D42.

Henderson:2023:CLM

[329] Leah Henderson. Crossing levels: Meta-induction and the problem of induction. Philosophy of Science, 90(4):985–993, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/crossing-levels-metainduction-
and-the-problem-of-induction/07085A3A1E810B4A95FC6BA80E45586D.

Dresow:2023:AUA

[330] Max Dresow. Actualism and uniformitarianism: From abstract commitments to forms of practice. Philosophy of Science, 90(4):994–1000, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/actualism-and-uniformitarianism-
from-abstract-commitments-to-forms-of-practice/49A787C396625BA009F36BD66CF49414.

Devanesan:2023:RCP

[331] Arjun Devanesan. The reference class problem and probabilities in the individual case: a response to Fuller. Philosophy of Science, 90 (4):1001–1009, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/reference-class-
problem-and-probabilities-in-the-individual-case-a-response-
to-fuller/5A099675AB3303A807C739D0D01F75B5.

Crull:2023:RKK

[332] Elise Crull. Review of Konrad Kleinknecht’s Werner Heisenberg: Reality and Its Order — Konrad Kleinknecht, ed., Werner Heisenberg: Reality and Its Order, with introduction by H. Rechenberg and commentary by E. P. Fischer, translated by M. B. Rumscheidt, N. Lukens, and I. Heisenberg. Cham, Switzerland: Springer (2019), 148 pp., $129.99. Philosophy of Science, 90(4):1010–1013, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-konrad-kleinknechts-
werner-heisenberg-reality-and-its-order-konrad-kleinknecht-ed-
werner-heisenberg-reality-and-its-order-with-introduction-by-
h-rechenberg-and-commentary-by-e-p-fischer-translated-by-m-b-
rumscheidt-n-lukens-and-i-heisenberg-cham-switzerland-springer-
2019-148-pp-12999/0C9AC6F396502302FB501A1DE763F381. See corri-

gendum [519].

Dong:2023:RJM

[333] Hao Dong. Review of Jeffrey McDonough’s A Miracle Creed: The Principle of Optimality in Leibniz’s Physics and Philosophy — Jeffrey Mc-Donough, A Miracle Creed: The Principle of Optimality in Leibniz’s Physics and Philosophy. Oxford: Oxford University Press (2022), 234 pp. $74.00 (hardcover). Philosophy of Science, 90(4):1013–1016, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-jeffrey-mcdonoughs-a-miracle-creed-
the-principle-of-optimality-in-leibnizs-physics-and-philosophy-
jeffrey-mcdonough-a-miracle-creed-the-principle-of-optimality-
in-leibnizs-physics-and-philosophy-oxford-oxford-university-
press-2022-234-pp-7400-hardcover/174DF07156F2D64A21AAD5B4B77DACC1.

Anonymous:2023:PSN

[334] Anonymous. Philosophy of Science 89, no. 5 cover and front matter — erratum. Philosophy of Science, 90(4):1017, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-

science/article/philosophy-of-science-89-no-5-cover-and-front-
matter-erratum/693C60B4BBE2A7EE670414D6E9E09756.

Anonymous:2023:PVIg

[335] Anonymous. PSA volume 90 issue 4 cover and front matter. Philosophy of Science, 90(4):f1–f4, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
4-cover-and-front-matter/5D6EF91DF81F3A8D5D526B8001AD6ABB.

Anonymous:2023:PVIh

[336] Anonymous. PSA volume 90 issue 4 cover and back matter. Philosophy of Science, 90(4):b1–b2, October 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
4-cover-and-back-matter/90B79C012B06B1F5BC13E50E5FCFD253.

Andersen:2023:P

[337] Holly K. Andersen. Preface. Philosophy of Science, 90(5):1019–1020, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/preface/0D794E5E7110236E63BDFEB15E0797ED.

Novick:2023:NTC

[338] Rose Novick. The neutral theory of conceptual complexity. Philosophy of Science, 90(5):1021–1030, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/neutral-theory-of-
conceptual-complexity/3A616D9B882A550750F78247F2C501BD.

Magnus:2023:SON

[339] P. D. Magnus. Scurvy and the ontology of natural kinds. Philosophy of Science, 90(5):1031–1039, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/scurvy-and-the-ontology-
of-natural-kinds/6E490FEC9C3151F502DB6ACC47903954.

Mostajir:2023:CAP

[340] Parysa Clare Mostajir. Classical American pragmatism as antiscientism. Philosophy of Science, 90(5):1040–1049, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/classical-american-pragmatism-as-antiscientism/
A58772CE8075E136D6720346463C4D69.

Fischer:2023:NFL

[341] Enno Fischer. Naturalness and the forward-looking justification of scientific principles. Philosophy of Science, 90(5):1050–1059, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/naturalness-and-the-forwardlooking-justification-
of-scientific-principles/950D8C5847ECFD5E11B1A957A77BF90F.

Springle:2023:TTM

[342] Alison Springle, Rebecca Dreier, and Seth Goldwasser. Trusting traumatic memory: Considerations from memory science. Philosophy of Science, 90(5):1060–1068, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/trusting-traumatic-
memory-considerations-from-memory-science/C87C5B3D4CA6D451194ACCC4A1187951.

Levy:2023:CBM

[343] Arnon Levy. Can Bayesian models of cognition show that we are (epistemically) rational? Philosophy of Science, 90(5):1069–1078, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/can-bayesian-models-of-cognition-show-that-we-
are-epistemically-rational/F5693B6DC765875A6AED6F285488D27D.

Curtis-Trudel:2023:MEC

[344] André Curtis-Trudel. Mathematical explanation in computer science. Philosophy of Science, 90(5):1079–1088, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/mathematical-explanation-in-computer-science/
2248FFA32F80510EAA9664B8A90CDD37.

Duede:2023:DLO

[345] Eamon Duede. Deep learning opacity in scientific discovery. Philosophy of Science, 90(5):1089–1099, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/deep-learning-opacity-
in-scientific-discovery/C46306D902A7AC87FD192D996639784A.

DiMarco:2023:CET

[346] Marina DiMarco. Cooperative epistemic trustworthiness. Philosophy of Science, 90(5):1100–1109, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/cooperative-
epistemic-trustworthiness/8105448C9862B891E717E8752DB33439.

Yee:2023:IDD

[347] Adrian K. Yee. Information deprivation and democratic engagement. Philosophy of Science, 90(5):1110–1119, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/information-deprivation-and-democratic-engagement/
E4447DCB5711FF217233D6F63FEF64D0.

MacLeod:2023:PMS

[348] Miles MacLeod and Michiru Nagatsu. Participatory modeling in sustainability science: The road to value-neutrality. Philosophy of Science, 90(5): 1120–1129, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/participatory-modeling-
in-sustainability-science-the-road-to-valueneutrality/63DD264E902C4CFCE4AAF7D659D06A32.

Belot:2023:DCD

[349] Gordon Belot. That does not compute: David Lewis on credence and chance. Philosophy of Science, 90(5):1130–1139, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/that-does-not-compute-david-lewis-on-credence-
and-chance/A564C0A6625EC6BDDA4B70AB96DDF976.

Kostic:2023:URG

[350] Milana Kostić. On the utility of research into geoengineering technologies for risk-avoidant agents. Philosophy of Science, 90(5):1140–1149, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/on-the-utility-of-research-into-
geoengineering-technologies-for-riskavoidant-agents/1224223F3CD1584944573D8AF56D25CC.

Schupbach:2023:LSB

[351] Jonah N. Schupbach. On the logical structure of best explanations. Philosophy of Science, 90(5):1150–1160, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/on-the-logical-structure-of-best-explanations/
7A490FC9EDA7617E76F025CC76312321.

Kinney:2023:CHS

[352] David Kinney. Causal history, statistical relevance, and explanatory power. Philosophy of Science, 90(5):1161–1172, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/causal-history-statistical-relevance-and-explanatory-
power/57625834E512134B882653EE4B3036F6.

Vieland:2023:AZE

[353] Veronica J. Vieland and Sang-Cheol Seok. Absolutely zero evidence. Philosophy of Science, 90(5):1173–1182, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/absolutely-zero-evidence/710AD5FB5F8405D62AC3675D2F2DBDA7.

Jantgen:2023:HME

[354] Ina Jäntgen. How to measure effect sizes for rational decision making. Philosophy of Science, 90(5):1183–1193, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/how-to-measure-effect-sizes-for-rational-decision-
making/E93D0C48D4C6B14E2E4A990530F84580.

Knox:2023:IDP

[355] Bennett Knox. The institutional definition of psychiatric condition and the role of well-being in psychiatry. Philosophy of Science, 90(5):1194– 1203, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/institutional-definition-of-psychiatric-
condition-and-the-role-of-wellbeing-in-psychiatry/B6679DD54B0CEAFEE169FA5EDE7628A6.

Erasmus:2023:BDM

[356] Adrian Erasmus. The bias dynamics model: Correcting for meta-biases in therapeutic prediction. Philosophy of Science, 90(5):1204–1213, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/bias-dynamics-model-correcting-for-metabiases-
in-therapeutic-prediction/BEE84865C520DA6DEA1A57C273E011C8.

Goldwasser:2023:FNA

[357] Seth Goldwasser. Finding normality in abnormality: On the ascription of normal functions to cancer. Philosophy of Science, 90(5):1214– 1223, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/finding-normality-in-
abnormality-on-the-ascription-of-normal-functions-to-cancer/
BA6C6DC1EEADC5047CD329DBBB3F2976.

Neto:2023:SCC

[358] Celso Neto and Letitia Meynell. Scaffold: a causal concept for evolutionary explanations. Philosophy of Science, 90(5):1224–1233, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/scaffold-a-causal-concept-for-
evolutionary-explanations/E4C8CCFD43E98509365EBCA336EB461E.

DiFrisco:2023:ATI

[359] James DiFrisco and Grant Ramsey. Adaptationism and trait individuation. Philosophy of Science, 90(5):1234–1243, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/adaptationism-and-trait-individuation/CDBD53CBCDA5B354A188D880FEF7831D.

Abrams:2023:RFP

[360] Marshall Abrams. Random foraging and perceived randomness. Philosophy of Science, 90(5):1244–1254, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/random-
foraging-and-perceived-randomness/163C7A034D4FA16511D7B9398A9D2DAD.

Browning:2023:VIS

[361] Heather Browning. Validating indicators of subjective animal welfare. Philosophy of Science, 90(5):1255–1264, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/validating-indicators-of-subjective-animal-welfare/
59501C4EB8E6E24A33D103FF67CD3C32.

Fagan:2023:IPH

[362] Melinda Bonnie Fagan. Immunology and population health: Collaboration without convergence. Philosophy of Science, 90(5):1265–1275, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/immunology-and-population-health-collaboration-
without-convergence/1DDBEB337E51437CCE3E9BCC02F70DE7.

Sung:2023:AEP

[363] Doohyun Sung and Bennett Holman. Against evidential pluralism in pharmaceutical regulation. Philosophy of Science, 90(5):1276–1285, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/against-evidential-pluralism-in-pharmaceutical-
regulation/ADB4B90B3EC8A32C80C53BF4DFCB91C4.

Elliott:2023:IVI

[364] Steve Elliott, Gina Eosco, Laura Newcomb, and Joseph Conran. Institutional values influence the design and evaluation of transition knowledge in funding proposals at NOAA. Philosophy of Science, 90(5):1286– 1296, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/institutional-values-
influence-the-design-and-evaluation-of-transition-knowledge-
in-funding-proposals-at-noaa/A34E0B2B95582BCEB4B7D668439BCAC8.

Shaw:2023:PRI

[365] Jamie Shaw. Peer review, innovation, and predicting the future of science: The scope of lotteries in science funding policy. Philosophy of Science, 90(5):1297–1306, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/peer-review-
innovation-and-predicting-the-future-of-science-the-scope-of-
lotteries-in-science-funding-policy/780541A102BC4DF6FA2F87BDF0FAC256.

Chua:2023:FAS

[366] Eugene Yew Siang Chua. T falls apart: On the status of classical temperature in relativity. Philosophy of Science, 90(5):1307– 1319, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/t-falls-apart-on-the-
status-of-classical-temperature-in-relativity/43CE3D77A75E3BF5DE46BF37F5671AE1.

Robertson:2023:TS

[367] Katie Robertson and Carina Prunkl. Is thermodynamics subjective? Philosophy of Science, 90(5):1320–1330, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/is-thermodynamics-subjective/11A3089569F8116868540104A9B63F19.

Manchak:2023:SSI

[368] Jb Manchak. On the (in?)Stability of spacetime inextendibility. Philosophy of Science, 90(5):1331–1341, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/on-
the-instability-of-spacetime-inextendibility/64967339482E13BAD36CD00837B57A3E.

Weatherall:2023:WDG

[369] James Owen Weatherall. Where does general relativity break down? Philosophy of Science, 90(5):1342–1351, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/where-
does-general-relativity-break-down/29A158BBF428113F6AB9F2AFF5B81DCE.

Jacobs:2023:DSD

[370] Caspar Jacobs. Are dynamic shifts dynamical symmetries? Philosophy of Science, 90(5):1352–1362, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/are-dynamic-
shifts-dynamical-symmetries/A5E60F5A766B87F4E555726416E0D18E.

Linnemann:2023:GCS

[371] Niels Linnemann, Chris Smeenk, and Mark Robert Baker. GR as a classical spin-2 theory? Philosophy of Science, 90(5):1363–1373, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/gr-as-a-classical-spin2-theory/9B0C95F3AFD5043676AEED650E8066EF.

Bursten:2023:MMN

[372] Julia R. S. Bursten and Kelle Dhein. Multiscale modeling in neuroethology: The significance of the mesoscale. Philosophy of Science, 90(5):1374– 1384, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/multiscale-modeling-in-neuroethology-
the-significance-of-the-mesoscale/593383A6ADA2AFC393413B91B91D004B.

Yao:2023:ESH

[373] Siyu Yao. Excavation in the sky: Historical inference in astronomy. Philosophy of Science, 90(5):1385–1395, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/excavation-in-the-sky-historical-inference-in-
astronomy/42C875A0394B1CE584F6ABD9266C4730.

Ohnesorge:2023:EPM

[374] Miguel Ohnesorge. The epistemic privilege of measurement: Motivating a functionalist account. Philosophy of Science, 90(5):1396–1406, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/epistemic-privilege-of-measurement-motivating-
a-functionalist-account/90BD2244F61A451963165A2771890A07.

Gluck:2023:RCM

[375] Stuart Gluck. Robustness of climate models. Philosophy of Science, 90(5): 1407–1416, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/robustness-of-climate-
models/1DA4E29B94254408B27328541E981E97.

Dethier:2023:API

[376] Corey Dethier. Against “Possibilist” interpretations of climate models. Philosophy of Science, 90(5):1417–1426, December 2023. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/against-possibilist-interpretations-of-climate-
models/B9F3F3403C747FDC6C55D335BDB74C82.

Feiten:2023:MTR

[377] Tim Elmo Feiten. The map/territory relationship in game-theoretic modeling of cultural evolution. Philosophy of Science, 90(5):1427–1436, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/mapterritory-relationship-in-gametheoretic-
modeling-of-cultural-evolution/4B96728A77C69E15DDC69A5959D5475F.

Heydon:2023:EEP

[378] Emily Heydon. Exploring an evolutionary paradox: an analysis of the “Spite Effect” and the “Nearly Neutral Effect” in synergistic models of finite populations. Philosophy of Science, 90(5):1437–1448, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/exploring-an-evolutionary-paradox-an-analysis-
of-the-spite-effect-and-the-nearly-neutral-effect-in-synergistic-
models-of-finite-populations/70E83190FAA38FF1983F8C3E464D3F45.

Kuorikoski:2023:EVM

[379] Jaakko Kuorikoski and Caterina Marchionni. Evidential variety and mixed-methods research in social science. Philosophy of Science, 90(5): 1449–1458, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/evidential-variety-and-
mixedmethods-research-in-social-science/3C33467351B9B26DE7836A0FFDBAA6F4.

vanBasshuysen:2023:AME

[380] Philippe van Basshuysen. Austinian model evaluation. Philosophy of Science, 90(5):1459–1468, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/austinian-model-
evaluation/F8DA9F6C5CFD1C1E1059A32D15EB023F.

Callender:2023:TNI

[381] Craig Callender. Temporal neutrality implies exponential temporal discounting. Philosophy of Science, 90(5):1469–1479, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/temporal-neutrality-implies-exponential-temporal-
discounting/5BDB1A33193E201D063162FFDDCB49EE.

Wylie:2023:PFF

[382] Alison Wylie. Philosophy of the field, in the field: Philosophy of Science Association 2020/2021 Presidential address. Philosophy of Science, 90(5): 1480–1495, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/philosophy-of-the-field-
in-the-field-philosophy-of-science-association-20202021-presidential-
address/684F1FAF0BE38C8F7E16AA9EC94CF460.

Anonymous:2023:PVIi

[383] Anonymous. PSA volume 90 issue 5 cover and front matter. Philosophy of Science, 90(5):f1–f7, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
5-cover-and-front-matter/4467246EB8242CD7EDE5C437D203008D.

Anonymous:2023:PVIj

[384] Anonymous. PSA volume 90 issue 5 cover and back matter. Philosophy of Science, 90(5):b1–b2, December 2023. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-90-issue-
5-cover-and-back-matter/533F500077ADDC8BA7C178265F27C0EA.

Eva:2024:CL

[385] Benjamin Eva. Comparative learning. Philosophy of Science, 91(1): 1–17, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/comparative-learning/
2336882FBAC6AB00E3AB213B4CF43A7B.

Parker:2024:EPA

[386] Wendy S. Parker. The epistemic projection approach to values in science. Philosophy of Science, 91(1):18–36, January 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/epistemic-projection-approach-to-values-in-science/
E8F846E15384A90AC6E15FA03BD41B4F.

Evans:2024:HEQ

[387] Peter W. Evans, Dominik Hangleiter, and Karim P. Y. Thébault. How to engineer a quantum wavefunction. Philosophy of Science, 91(1):37–57, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/how-to-engineer-
a-quantum-wavefunction/E5E101488CD831BA3C64A9F80173F5E1.

Hundertmark:2024:DAT

[388] Fabian Hundertmark. The dilemma of ahistorical teleosemantics. Philosophy of Science, 91(1):58–71, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/dilemma-of-ahistorical-
teleosemantics/8B3A2F860E53464F71785D710E4E36FB.

DiMarco:2024:MD

[389] Marina DiMarco. Models as dogwhistles. Philosophy of Science, 91(1): 72–89, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/models-as-dogwhistles/86056B9BD5627BAAD25D1824C431046E.

Brown:2024:EMU

[390] Simon Alexander Burns Brown. Episodic memory and unrestricted learning. Philosophy of Science, 91(1):90–110, January 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/episodic-memory-and-unrestricted-learning/E769ED00F33D76545BE6B1C9A89FBA3D.

Boyd:2024:OEA

[391] Nora Mills Boyd and Dana Matthiessen. Observations, experiments, and arguments for epistemic superiority in scientific methodology. Philosophy of Science, 91(1):111–131, January 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/observations-experiments-and-arguments-for-epistemic-
superiority-in-scientific-methodology/F5AC61A0079256D8E4109665D3B8B595.

Siegel:2024:PLM

[392] Gabriel Siegel and Carl F. Craver. Phenomenological laws and mechanistic explanations. Philosophy of Science, 91(1):132–150, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/phenomenological-laws-and-mechanistic-explanations/
2D34AA6420210D1C44E175C21B5D3BAD.

Malaterre:2024:EMS

[393] Christophe Malaterre and Martin Léonard. Epistemic markers in the scientific discourse. Philosophy of Science, 91(1):151–174, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/epistemic-markers-in-the-scientific-discourse/
E256184D590F833A2255790CA9DAFD1A.

Babic:2024:MRS

[394] Joshua Babic and Lorenzo Cocco. Mandersian relationism: Space, modality, and equivalence. Philosophy of Science, 91(1):175–185, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/mandersian-relationism-space-modality-and-equivalence/
D7808D9A9C2CFB479437D67F047A44D7.

Zollman:2024:AJI

[395] Kevin J. S. Zollman, Julian García, and Toby Handfield. Academic journals, incentives, and the quality of peer review: a model. Philosophy of Science, 91(1):186–203, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/academic-journals-
incentives-and-the-quality-of-peer-review-a-model/04F05FC5C66BF9B8BDE8CAD607BF1933.

Thoma:2024:MMP

[396] Johanna Thoma. Merely means paternalist? Prospect theory and “Debiased” welfare analysis. Philosophy of Science, 91(1):204–224, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/merely-means-paternalist-prospect-theory-and-
debiased-welfare-analysis/DF06037C2D7F741CB3B9A333155B58A4.

Schneider:2024:RFS

[397] Mike D. Schneider. A role for the Fauxrizon in the semiclassical limit of a fuzzball. Philosophy of Science, 91(1):225–242, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/role-for-the-fauxrizon-in-the-semiclassical-
limit-of-a-fuzzball/DDDC602CA7803DC35BF70EAF4B1EC145.

Pregel:2024:RSN

[398] Fabian Pregel. Reply to Sprenger’s “A Novel Solution to the Problem of Old Evidence”. Philosophy of Science, 91(1):243–252, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/reply-to-sprengers-a-novel-solution-to-the-problem-
of-old-evidence/941322E7A270F11FD97296BFB06C0181.

Westerblad:2024:RMM

[399] Oscar Westerblad. Review of Michela Massimi and Casey D. Mc-Coy’s Understanding Perspectivism: Scientific Challenges and Methodological Prospects — Michela Massimi and Casey D. McCoy, Understanding Perspectivism: Scientific Challenges and Methodological Prospects. New York: Routledge (2020), 210 pp. $136 (hardcover). Philosophy of Science, 91(1):253–257, January 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-michela-massimi-and-casey-d-mccoys-
understanding-perspectivism-scientific-challenges-and-methodological-
prospects-michela-massimi-and-casey-d-mccoy-understanding-perspectivism-
scientific-challenges-and-methodological-prospects-new-york-
routledge-2020-210-pp-136-hardcover/7E75998213A9649B5FDFF93EED0EB058.

LaCroix:2024:RRJ

[400] Travis LaCroix. Review of Ronald J. Planer and Kim Sterelny’s From Signal to Symbol: The Evolution of Language — Ronald J.

Planer and Kim Sterelny, From Signal to Symbol: The Evolution of Language. Cambridge: The MIT Press (2021), 296 pp. $35.00 (hardcover). Philosophy of Science, 91(1):257–260, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-ronald-j-planer-and-kim-sterelnys-
from-signal-to-symbol-the-evolution-of-language-ronald-j-planer-
and-kim-sterelny-from-signal-to-symbol-the-evolution-of-language-
cambridge-the-mit-press-2021-296-pp-3500-hardcover/88486B72139EEA3FD24D4BC7640277D9.

Nyrup:2024:RER

[401] Rune Nyrup. Review of Emanuele Ratti and Thomas A. Stapleford’s Science, Technology, and Virtues: Contemporary Perspectives — Emanuele Ratti and Thomas A. Stapleford, eds., Science, Technology, and Virtues: Contemporary Perspectives. Oxford: Oxford University Press (2021), 312 pp. $99 (hardcover). Philosophy of Science, 91(1):261–263, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-emanuele-ratti-and-thomas-a-staplefords-
science-technology-and-virtues-contemporary-perspectives-emanuele-
ratti-and-thomas-a-stapleford-eds-science-technology-and-virtues-
contemporary-perspectives-oxford-oxford-university-press-2021-
312-pp-99-hardcover/1F5B94658A38F27327FF19D5997E6FC3.

Shech:2024:RRW

[402] Elay Shech. Review of Robert W. Batterman’s A Middle Way: a Non-Fundamental Approach to Many-Body Physics — Robert W. Batterman, A Middle Way: a Non-Fundamental Approach to Many-Body Physics. New York: Oxford University Press (2021), 174 pp. $74 (hardcover). Philosophy of Science, 91(1):264–268, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-robert-w-battermans-
a-middle-way-a-nonfundamental-approach-to-manybody-physics-robert-
w-batterman-a-middle-way-a-nonfundamental-approach-to-manybody-
physics-new-york-oxford-university-press-2021-174-pp-74-hardcover/
83FFEB98C61475A1F0918E664240318B.

Anonymous:2024:PVIa

[403] Anonymous. PSA volume 91 issue 1 cover and front matter. Philosophy of Science, 91(1):f1–f5, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
1-cover-and-front-matter/508E9CD91219EAFFC1F91AC923FE88EA.

Anonymous:2024:PVIb

[404] Anonymous. PSA volume 91 issue 1 cover and back matter. Philosophy of Science, 91(1):b1–b2, January 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
1-cover-and-back-matter/FFBABDD7EABE90719E422A2812092719.

Vredenburgh:2024:CER

[405] Kate Vredenburgh. Causal explanation and revealed preferences. Philosophy of Science, 91(2):269–287, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/causal-explanation-
and-revealed-preferences/C8980E6203605D236C4BF489690CAE58.

Franklin:2024:INJ

[406] Alexander Franklin. Incoherent? No, just decoherent: How quantum many worlds emerge. Philosophy of Science, 91(2):288–309, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/incoherent-no-just-decoherent-
how-quantum-many-worlds-emerge/3DF319C9E1FF87BD5EADE05C9EF350E8.

Evron:2024:WDS

[407] Aya Evron. What do sexes have to do with (models of) sexual selection? Philosophy of Science, 91(2):310–328, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/what-do-sexes-have-to-do-with-models-of-sexual-
selection/BE4FA2DA65882483AE333088C2472450.

Gabriel:2024:CCB

[408] Nathan Gabriel and Cailin O’Connor. Can confirmation bias improve group learning? Philosophy of Science, 91(2):329–350, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/can-confirmation-bias-improve-group-learning/
8467624293085CF214CB2D20AB4CFFAC.

Veldman:2024:MMC

[409] Michael Veldman. Mathematizing metaphysics: The case of the principle of least action. Philosophy of Science, 91(2):351–369, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/mathematizing-metaphysics-the-
case-of-the-principle-of-least-action/A6CFB59070F4115CE50CE3C092048177.

Artiga:2024:DM

[410] Marc Artiga and Cédric Paternotte. Deception as mimicry. Philosophy of Science, 91(2):370–389, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/deception-as-mimicry/
0D960CBFB788E628C7EDB0C9CD9B4C69.

Lavelle:2024:GUU

[411] Jane Suilin Lavelle. Growth from uncertainty: Understanding the replication ‘crisis’ in infant cognition. Philosophy of Science, 91(2):390–409, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/growth-from-uncertainty-understanding-the-replication-
crisis-in-infant-cognition/A432336A2364A9F8954A0765CE5CC155.

Dorst:2024:DBS

[412] Chris Dorst. Does the best system need the past hypothesis? Philosophy of Science, 91(2):410–429, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/does-the-best-system-
need-the-past-hypothesis/8283DFB5F076E6D445DC61B0D1DE87CF.

Sims:2024:PDH

[413] Matthew Sims. The principle of dynamic holism: Guiding methodology for investigating cognition in nonneuronal organisms. Philosophy of Science, 91(2):430–448, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/principle-of-dynamic-
holism-guiding-methodology-for-investigating-cognition-in-nonneuronal-
organisms/2B21BE9620E4AFB00579298709CB9209.

Thorell:2024:HDB

[414] Amanda Thorell. Health and disease: Between naturalism and normativism. Philosophy of Science, 91(2):449–467, April 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/health-and-disease-between-naturalism-and-normativism/
68679A4BFB6C798367E69004A953E3D9.

Verhaegh:2024:RRA

[415] Sander Verhaegh. The reception of relativity in American philosophy. Philosophy of Science, 91(2):468–487, April 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/reception-of-relativity-in-american-philosophy/
F3B4016B85157C266A8390DC95717704.

Tasdan:2024:SCR

[416] Ufuk I. Tasdan and Karim P. Y. Thébault. Spacetime conventionalism revisited. Philosophy of Science, 91(2):488–509, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/spacetime-conventionalism-revisited/105E7E6F88395608E0E8AA87F9C50766.

Kurth:2024:EAG

[417] Charlie Kurth. An evolutionary account of guilt? Philosophy of Science, 91(2):510–518, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/an-evolutionary-
account-of-guilt/30B24ACB7E2795043F56E37B3BACDB17.

Howard:2024:BRM

[418] Don Howard. Book review: Matthias Neuber and Adam Tamas Tuboly, eds., Ernest Nagel: Philosophy of Science and the Fight for Clarity. Cham, Switzerland: Springer. 348 pp. $159.99. Hardcover and softcover (2022). Philosophy of Science, 91(2):519–521, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/matthias-neuber-and-adam-tamas-tuboly-eds-ernest-
nagel-philosophy-of-science-and-the-fight-for-clarity-cham-switzerland-
springer-348-pp-15999-hardcover-and-softcover-2022/A3EFB88E8DDFDB0C7688E48E93D22E8E.

Carrera:2024:RCD

[419] Matilde Carrera and Aja Watkins. Review of caitlin donahue Wylie’s preparing dinosaurs: The work behind the scenes — caitlin donahue wylie, preparing dinosaurs: The work behind the scenes. Cambridge, MA: MIT Press (2021), 264 pp. $75.00 (paperback). Philosophy of Science, 91(2): 522–525, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-caitlin-donahue-wylies-
preparing-dinosaurs-the-work-behind-the-scenes-caitlin-donahue-

wylie-preparing-dinosaurs-the-work-behind-the-scenes-cambridge-
ma-mit-press-2021-264-pp-7500-paperback/FB3C3C50543EF3C7D1FBEEF471235688.

Cuffaro:2024:RSP

[420] Michael E. Cuffaro. Review of Slobodan Perović’s From Data to Quanta — Niels Bohr’s Vision of Physics. Chicago, IL: The University of Chicago Press (2021), 280 pp., $45 (cloth). Philosophy of Science, 91(2):525– 529, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-slobodan-perovics-from-
data-to-quanta-niels-bohrs-vision-of-physics-slobodan-perovi-
c-from-data-to-quanta-niels-bohrs-vision-of-physics-chicago-
il-the-university-of-chicago-press-2021-280-pp-45-cloth/F70FD92FEA6B1ECC0A3F80FFE82A8124.

Anonymous:2024:PVIc

[421] Anonymous. PSA volume 91 issue 2 cover and front matter. Philosophy of Science, 91(2):f1–f4, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
2-cover-and-front-matter/721D9EBCD9986AC2FD27AC214667DB38.

Anonymous:2024:PVId

[422] Anonymous. PSA volume 91 issue 2 cover and back matter. Philosophy of Science, 91(2):b1–b2, April 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
2-cover-and-back-matter/1FFD12D8D67AF20132BC3D3225046577.

Urbach:2024:MCH

[423] Peter Urbach. In memoriam: Colin Howson. Philosophy of Science, 91(3):531–532, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/in-memoriam-
colin-howson/9318D7CEF2AA283DAA5F03AD3465120A.

Howson:2024:BFA

[424] Colin Howson. Beyond finite additivity. Philosophy of Science, 91(3): 533–542, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/beyond-finite-additivity/696609F39D3AD96AEEFB3941EC5978F9.

Stegenga:2024:JSP

[425] Jacob Stegenga. Justifying scientific progress. Philosophy of Science, 91 (3):543–560, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/justifying-scientific-progress/
4A66629E53CFA195D11CB6E9D7B918AB.

Griffith:2024:PAP

[426] Trevor Griffith and Adrian Kind. Pain asymbolia is not pain. Philosophy of Science, 91(3):561–578, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/pain-asymbolia-is-
not-pain/14D42EF12870DC9C6BA506F2C5F5AC5F.

Huang:2024:LBU

[427] Alice C. W. Huang. Landscapes and bandits: a unified model of functional and demographic diversity. Philosophy of Science, 91(3):579–594, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/landscapes-and-bandits-a-unified-model-of-functional-
and-demographic-diversity/7D6ADD1049AF38C3F34AC2DF64A76FF1.

Hazelwood:2024:EDR

[428] Caleb Hazelwood. An emerging dilemma for reciprocal causation. Philosophy of Science, 91(3):595–614, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/an-emerging-
dilemma-for-reciprocal-causation/BDC1761CDCE71DEE349F4026FF8F10E5.

Gamboa:2024:CMO

[429] J. P. Gamboa. On cognitive modeling and other minds. Philosophy of Science, 91(3):615–633, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/on-cognitive-modeling-
and-other-minds/00F301790F52C52D6091163020F79281.

Blando:2024:PP

[430] Francesca Zaffora Blando. Pride and probability. Philosophy of Science, 91 (3):634–660, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/pride-and-probability/CB15DB2546E744E9DCDDB2F4D14633B4.

Peters:2024:HGP

[431] Uwe Peters and Olivier Lemeire. Hasty generalizations are pervasive in experimental philosophy: a systematic analysis. Philosophy of Science, 91 (3):661–681, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/hasty-generalizations-are-pervasive-
in-experimental-philosophy-a-systematic-analysis/27AF62B58334B5120F3317CE9E6259C7.

Mollo:2024:FCR

[432] Dimitri Coelho Mollo and Alfredo Vernazzani. The formats of cognitive representation: a computational account. Philosophy of Science, 91(3):682–701, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/formats-of-cognitive-
representation-a-computational-account/DAA745855F0754879B4D83DC698B894A.

Zanetti:2024:CPR

[433] Luca Zanetti. Confidence in probabilistic risk assessment. Philosophy of Science, 91(3):702–720, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/confidence-in-probabilistic-
risk-assessment/E7F3CBDE94325C92D5DBFE30A5FAD235.

Mancosu:2024:TRC

[434] Paolo Mancosu and Guillaume Massas. Totality, regularity, and cardinality in probability theory. Philosophy of Science, 91(3):721–740, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/totality-regularity-and-cardinality-in-probability-
theory/1FABB0AD8522D2634335CC584D58BF56.

Benedetto:2024:TCN

[435] Matteo De Benedetto and Michele Luchetti. Theory choice as niche construction: The feedback loop between scientific theories and epistemic values. Philosophy of Science, 91(3):741–758, July 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/theory-choice-as-niche-construction-the-feedback-
loop-between-scientific-theories-and-epistemic-values/1A42B6E147D2009CC7C3B781377FC9FA.

Wang:2024:CBD

[436] Minkyung Wang and Chisu Kim. Credence and belief: Distance- and utility-based approaches. Philosophy of Science, 91(3):759–779, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-

science/article/credence-and-belief-distance-and-utilitybased-
approaches/04B64AAB64C69F63377FCACB20A42847.

Schurz:2024:HMI

[437] Gerhard Schurz. Henderson on meta-induction and the problem of induction. Philosophy of Science, 91(3):780–785, July 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/henderson-on-metainduction-and-the-problem-of-
induction/323DB81C8D244F90B832742F17FC4D58.

Povich:2024:RSI

[438] Mark Povich. Review of Stavros Ioannidis and Stathis Psillos’s Mechanisms in Science: Method or Metaphysics? — Stavros Ioannidis and Stathis Psillos, Mechanisms in Science: Method or Metaphysics? Cambridge: Cambridge University Press (2022), 250 pp. $99.99 (hardback). Philosophy of Science, 91(3):786–789, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-stavros-ioannidis-and-stathis-psilloss-
mechanisms-in-science-method-or-metaphysics-stavros-ioannidisand-
stathis-psillos-mechanisms-in-science-method-or-metaphysics-
cambridge-cambridge-university-press-2022-250-pp-9999-hardback/
B0F20EA85DF6EACBB8D8B928F1D50E95.

Bianchi:2024:RBA

[439] Silvia De Bianchi. Review of Boyd et al.’s Philosophy of Astrophysics: Stars, Simulations, and the Struggle to Determine What is Out There - “Philosophy of Astrophysics: Stars, Simulations, and the Struggle to Determine What is Out There”, edited by Nora Mills Boyd, Siska De Baerdemaeker, Kevin Heng, and Vera Matarese, Cham: Springer, 2023. Philosophy of Science, 91(3):789–792, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-boyd-et-als-philosophy-of-astrophysics-
stars-simulations-and-the-struggle-to-determine-what-is-out-
there-philosophy-of-astrophysics-stars-simulations-and-the-struggle-
to-determine-what-is-out-there-edited-by-nora-mills-boyd-siska-
de-baerdemaeker-kevin-heng-and-vera-matarese-cham-springer-2023/
0CB15D363C18229266F5E68C7C3CC8B7.

Anonymous:2024:PVIe

[440] Anonymous. PSA volume 91 issue 3 cover and front matter. Philosophy of Science, 91(3):f1–f4, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
3-cover-and-front-matter/BE4855554AC8E837B65C777878597A94.

Anonymous:2024:PVIf

[441] Anonymous. PSA volume 91 issue 3 cover and back matter. Philosophy of Science, 91(3):b1–b2, July 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
3-cover-and-back-matter/B78A003968512B92B93791376270F50F.

Watkins:2024:PPB

[442] Aja Watkins. Paleoclimate proxies and the benefits of disunity. Philosophy of Science, 91(4):793–810, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/paleoclimate-proxies-
and-the-benefits-of-disunity/E8444E45B7D218006C4CAD4938D8A5DA.

See erratum [442].

Schulz:2024:PEA

[443] Armin W. Schulz. Phylogenetic economics: Animal models and the study of choice. Philosophy of Science, 91(4):811–830, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/phylogenetic-economics-animal-
models-and-the-study-of-choice/EEAF6FEB7B4841203ED3F40538BDF113.

Ramirez:2024:SS

[444] Sebastián Murgueitio Ramírez. On symmetries and springs. Philosophy of Science, 91(4):831–850, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/on-symmetries-and-
springs/09DA7C816B63103418D9E5795E9EF582.

deMelo-Martin:2024:CAC

[445] Inmaculada de Melo-Martín. Concerns about contextual values in science and the legitimate/illegitimate distinction. Philosophy of Science, 91(4): 851–868, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/concerns-about-contextual-values-
in-science-and-the-legitimateillegitimate-distinction/701B6662937DAA22977F52BF89A962BF.

Ramsey:2024:DCE

[446] Grant Ramsey and Cristina Villegas. Developmental channeling and evolutionary dappling. Philosophy of Science, 91(4):869–886, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/developmental-channeling-and-evolutionary-dappling/
58C9D4C6DA1757199FEB672526492ED0.

Jha:2024:MEC

[447] Aditya Jha, Douglas Campbell, Clemency Montelle, and Phillip L. Wilson. Are mathematical explanations causal explanations in disguise? Philosophy of Science, 91(4):887–905, October 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/are-mathematical-explanations-causal-explanations-
in-disguise/DA2028D72FDB76C7275E4E215B9B5B51.

Liron:2024:TIC

[448] Amir Liron and Ittay Nissan-Rozen. Triangulation, incommensurability, and conditionalization. Philosophy of Science, 91(4):906–929, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/triangulation-incommensurability-and-conditionalization/
8A94F93F6DB7EBC6ED224920198D15B1.

Figdor:2024:ICC

[449] Carrie Figdor. Individuating cognitive characters: Lessons from praying mantises and plants. Philosophy of Science, 91(4):930–949, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/individuating-cognitive-characters-lessons-from-
praying-mantises-and-plants/73155E45022D59C8B04A5B16547821C4.

Garson:2024:FPS

[450] Justin Garson. Functions and populations: Sharpening the generalized selected effects theory of function. Philosophy of Science, 91(4):950– 968, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/functions-and-populations-sharpening-
the-generalized-selected-effects-theory-of-function/C6030AD49B5FDB1EBA4071281C80547D.

Guo:2024:TAR

[451] Bixin Guo. Two approaches to reduction: a case study from statistical mechanics. Philosophy of Science, 91(4):969–989, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/two-approaches-to-reduction-a-
case-study-from-statistical-mechanics/F59508B78D5BA5110A0EFD36296EFAA3.

Ward:2024:HHS

[452] Zina B. Ward and Kathleen A. Creel. To hedge or not to hedge: Scientific claims and public justification. Philosophy of Science, 91 (4):990–1008, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/to-hedge-or-not-
to-hedge-scientific-claims-and-public-justification/B58C1B12B2DB74225DBCCFAE2CCBB0F0.

Dicks:2024:BE

[453] Henry Dicks. Biomimetic epistemology. Philosophy of Science, 91(4):1009– 1029, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/biomimetic-epistemology/9F049DB2D88D32CF8DBA84DAEF751361.

Sterkenburg:2024:RJD

[454] Tom F. Sterkenburg. Review of John D. Norton’s The Material Theory of Induction — John D. Norton, The Material Theory of Induction. Calgary: University of Calgary Press (2021), 680 pp. $119.99 (hardcover). Philosophy of Science, 91(4):1030–1033, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-john-d-nortons-the-material-theory-
of-induction-john-d-norton-the-material-theory-of-induction-
calgary-university-of-calgary-press-2021-680-pp-11999-hardcover/
948C726BA0601ACBF87588FF0AF1998E.

Anonymous:2024:RPS

[455] Anonymous. Referees for Philosophy of Science. Philosophy of Science, 91 (4):1034–1040, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/referees-for-philosophy-
of-science/473DC4E31DD5865AD8BF311738CA20EE.

Anonymous:2024:PVIg

[456] Anonymous. PSA volume 91 issue 4 cover and front matter. Philosophy of Science, 91(4):f1–f4, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
4-cover-and-front-matter/467083E262D4D38675AB58DE631487EB.

Anonymous:2024:PVIh

[457] Anonymous. PSA volume 91 issue 4 cover and back matter. Philosophy of Science, 91(4):b1–b2, October 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
4-cover-and-back-matter/718A7A399A1FA908E5D0CCCF6BEF8AAE.

Andersen:2024:P

[458] Holly K. Andersen. Preface. Philosophy of Science, 91(5):1041–1042, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/preface/2158DD7FB39371B2058B84ADDA8AAE05.

Dupre:2024:DSU

[459] John Dupré. The disunity of science and the unity of the World Presidential Address, PSA 2022. Philosophy of Science, 91(5):1043–1057, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/disunity-of-science-and-the-unity-
of-the-world-presidential-address-psa-2022/6A53712FA32AB3C58A8D30F9BD303B64.

Goodwin:2024:SCI

[460] William Goodwin. Strategic constraints on interacting fields: Quantum chemistry and organic chemistry. Philosophy of Science, 91(5):1058– 1067, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/strategic-constraints-on-interacting-
fields-quantum-chemistry-and-organic-chemistry/F7EB029096FD335BF3F58150F0A9733B.

Ross:2024:CCL

[461] Lauren N. Ross. Causal constraints in the life and social sciences. Philosophy of Science, 91(5):1068–1077, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/causal-constraints-in-the-life-and-social-sciences/
7A928526E4FF4DC7E67B87487836B08A.

Kostic:2024:REC

[462] Daniel Kostić. On the role of erotetic constraints in noncausal explanations. Philosophy of Science, 91(5):1078–1088, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/on-the-role-of-erotetic-constraints-in-noncausal-
explanations/A6E6CCC04A493668B4180F574395ED17.

Weinberger:2024:ISD

[463] Naftali Weinberger. The insufficiency of statistics for detecting racial discrimination by police. Philosophy of Science, 91(5):1089–1097, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/insufficiency-of-statistics-for-detecting-racial-
discrimination-by-police/58938EE492F91AC6E286E704E50A6C11.

Tolbert:2024:CAA

[464] Alexander Williams Tolbert. Causal agnosticism about race: Variable selection problems in causal inference. Philosophy of Science, 91(5):1098– 1108, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/causal-agnosticism-about-race-
variable-selection-problems-in-causal-inference/EFF78D42C71F15B11C17EE043208A851.

Klein:2024:EAE

[465] Colin Klein. Exploratory analysis and the expected value of experimentation. Philosophy of Science, 91(5):1109–1117, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/exploratory-analysis-and-the-expected-value-
of-experimentation/73CC5FC7EB9520FDB907F4910C72A9FB.

Branch:2024:RCS

[466] T. Y. Branch and Heather Douglas. Rethinking the conceptual space for science in society after the VFI. Philosophy of Science, 91(5):1118– 1127, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/rethinking-the-conceptual-space-
for-science-in-society-after-the-vfi/E6462D298C1E6895C8C22932B8ED584C.

Bihan:2024:HSP

[467] Soazig Le Bihan. How to not secure public trust in science: Representative values versus polarization and marginalization. Philosophy of Science, 91(5):1128–1138, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/how-to-not-secure-
public-trust-in-science-representative-values-versus-polarization-
and-marginalization/9A7D89BC608414714D4B1A45B08D10CD.

Pinto:2024:PSC

[468] Manuela Fernández Pinto. Pandemic science and commercial values: An institutional account of values in science. Philosophy of Science, 91(5): 1139–1148, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/pandemic-science-and-
commercial-values-an-institutional-account-of-values-in-science/
64DC7D4021019C03BC7140FF4EE26333.

Douglas:2024:SIR

[469] Heather Douglas. Structuring institutions for responsible and accountable science. Philosophy of Science, 91(5):1149–1158, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/structuring-institutions-for-responsible-and-
accountable-science/8E8FBCDA377C118485FACC024D80E6C9.

Biddle:2024:OVS

[470] Justin B. Biddle. Organizations and values in science and technology. Philosophy of Science, 91(5):1159–1168, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/organizations-and-values-in-science-and-technology/
54CADA208B1DD089313F1C3AB71D8582.

Dang:2024:MRR

[471] Haixin Dang. Minority reports: Registering dissent in science. Philosophy of Science, 91(5):1169–1178, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/minority-
reports-registering-dissent-in-science/3750E24DFB4D34CED8E36507767B4D27.

Wong:2024:SDE

[472] Kekoa Wong and Hannah Rubin. Social dynamics and the evolution of disciplines. Philosophy of Science, 91(5):1179–1188, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/social-dynamics-and-the-evolution-of-disciplines/
D5F4AED5B4CDB764DAE6AB0D29CFF429.

Wu:2024:BTB

[473] Jingyi Wu. Better than best: Epistemic landscapes and diversity of practice in science. Philosophy of Science, 91(5):1189–1198, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/better-than-best-epistemic-landscapes-and-diversity-
of-practice-in-science/1125693C6EFCD39D13E3108840B87305.

Akagi:2024:SME

[474] Mikio Akagi. Structural microaggressions for explaining outcome gaps. Philosophy of Science, 91(5):1199–1209, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/structural-microaggressions-for-explaining-outcome-
gaps/940722948387B947802B61419E404797.

Parke:2024:ADS

[475] Emily C. Parke and Daniel Hikuroa. Against defending science: Asking better questions about indigenous knowledge and science. Philosophy of Science, 91(5):1210–1220, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/against-defending-
science-asking-better-questions-about-indigenous-knowledge-and-
science/4003F8B666972AE0CEF15F609A170810.

Ludwig:2024:TPS

[476] David Ludwig, Charbel N. El-Hani, Fabio Gatti, Catherine Kendig, Matthias Kramm, Lucia Neco, Abigail Nieves Delgado, Luana Poliseli, Vitor Renck, Adriana Ressiore C., Luis Reyes Galindo, Thomas Lloyd Rickard, Gabriela De La Rosa, Julia J. Turska, Francisco Vergara-Silva, and Robert A. Wilson. Transdisciplinary philosophy of science: Meeting the challenge of indigenous expertise. Philosophy of Science, 91(5):1221– 1231, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/transdisciplinary-philosophy-of-
science-meeting-the-challenge-of-indigenous-expertise/633BF57FC069DA7C472D07B705029224.

Smeenk:2024:DEM

[477] Chris Smeenk and James Owen Weatherall. Dark energy or modified gravity? Philosophy of Science, 91(5):1232–1241, December 2024.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/dark-energy-or-modified-gravity/96C7A6E56CAAE8A2BF96A72AABB6E929.

Dewar:2024:EC

[478] Neil Dewar. Equivalence and convention. Philosophy of Science, 91 (5):1242–1250, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/equivalence-and-
convention/50F1204948753928866661B5A31344FE.

Perry:2024:AQP

[479] Zee R. Perry. Against quantitative primitivism. Philosophy of Science, 91(5):1251–1261, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/against-quantitative-
primitivism/52805B49CCA4E66E14C1D678B55E7328.

Meskhidze:2024:TCS

[480] Helen Meskhidze and James Owen Weatherall. Torsion in the classical spacetime context. Philosophy of Science, 91(5):1262–1273, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/torsion-in-the-classical-spacetime-context/0AC21F88F934E0654AF565A64C6C8D97.

Bennett:2024:CRV

[481] Marissa Bennett and Michael E. Miller. The conventionality of real-valued quantities. Philosophy of Science, 91(5):1274–1284, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/conventionality-of-realvalued-
quantities/F5CA9978460723CE836BB80CFBE6CF7F.

Lin:2024:TNP

[482] Qiu Lin. Toward a new phase of Du Châtelet scholarship: From Institutions de Physique (1740) to Institutions Physiques (1742). Philosophy of Science, 91(5):1285–1295, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/toward-a-new-phase-
of-du-chatelet-scholarship-from-institutions-de-physique-1740-
to-institutions-physiques-1742/E2A64DE3E7EA3690242FC3C45CCAFD7B.

Clatterbuck:2024:HEG

[483] Hayley Clatterbuck. Hume’s externalist gambit. Philosophy of Science, 91(5):1296–1305, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/humes-externalist-
gambit/D6D2CF66B7E281E020AF4A054D673C6B.

Chakravartty:2024:RDA

[484] Anjan Chakravartty. Resolving debates about scientific realism: The challenge from stances. Philosophy of Science, 91(5):1306–1315, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/resolving-debates-about-scientific-realism-the-
challenge-from-stances/0D530375FEE14858467FF257EF602C7D.

Palluet:2024:ND

[485] Amélie Godard Palluet and Marie Gueguen. Navigating in the dark. Philosophy of Science, 91(5):1316–1326, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/navigating-in-the-dark/C02547A095F27160DE525E3B3FDA25CE.

Hoffman:2024:NHC

[486] Kate Nicole Hoffman and Karen Kovaka. A new heuristic for climate adaptation. Philosophy of Science, 91(5):1327–1337, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/new-heuristic-for-climate-adaptation/2D9A588F0A8258D15FD8821DB0C24AB5.

Leonelli:2024:DST

[487] Sabina Leonelli. Is data science transforming biomedical research? Evidence, expertise, and experiments in COVID-19 science. Philosophy of Science, 91(5):1338–1348, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/is-data-science-
transforming-biomedical-research-evidence-expertise-and-experiments-
in-covid19-science/3F5F54E46B2589537B38BDFDA2893D3E.

Chin-Yee:2024:CAP

[488] Benjamin Chin-Yee and Anya Plutynski. Concepts of actionability in precision oncology. Philosophy of Science, 91(5):1349–1360, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/concepts-of-actionability-in-precision-oncology/
1647196A542DBA7084C9B75BC61C1D30.

Feest:2024:WRC

[489] Uljana Feest. What is the replication crisis a crisis of? Philosophy of Science, 91(5):1361–1371, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/what-is-the-
replication-crisis-a-crisis-of/A9F78FD7A73226C9EB3580D3955108BA.

Cruwell:2024:PGT

[490] Sophia Crüwell. The psychologist’s green thumb. Philosophy of Science, 91(5):1372–1381, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psychologists-green-
thumb/A2A899C7F7276148575E98007A7C4313.

Colaco:2024:WMR

[491] David Colaço and Jonathan Najenson. Where memory resides: Is there a rivalry between molecular and synaptic models of memory? Philosophy of Science, 91(5):1382–1392, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/where-memory-resides-is-there-a-rivalry-between-
molecular-and-synaptic-models-of-memory/081F7702DAE6654F262B27F22F7C930C.

Doulcier:2024:ETI

[492] Guilhem Doulcier, Peter Takacs, and Pierrick Bourrat. Evolutionary transitions in individuality and life cycle closure. Philosophy of Science, 91(5): 1393–1403, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/evolutionary-transitions-
in-individuality-and-life-cycle-closure/D4B2EDAD4E9F1C831B1DE72429791B62.

Griesemer:2024:MSD

[493] James Griesemer. Modeling scaffolded development into the adjacent possible environment. Philosophy of Science, 91(5):1404–1414, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/modeling-scaffolded-development-into-the-adjacent-
possible-environment/40608856F38539A2E256CE65AD5FEA9C.

Millstein:2024:CDE

[494] Roberta L. Millstein. The coevolution of descriptive and evaluative beliefs in Aldo Leopold’s thinking. Philosophy of Science, 91 (5):1415–1424, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/coevolution-
of-descriptive-and-evaluative-beliefs-in-aldo-leopolds-thinking/
5C98AE5B0A2E0701C2658533043AE250.

Takacs:2024:FCT

[495] Peter Takacs, Guilhem Doulcier, and Pierrick Bourrat. From fitnesscentered to trait-centered explanations: What evolutionary transitions in individuality teach us about fitness. Philosophy of Science, 91(5): 1425–1434, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/from-fitnesscentered-
to-traitcentered-explanations-what-evolutionary-transitions-
in-individuality-teach-us-about-fitness/9BAD30D0D8BDD31237A6C7FB21323CB6.

Babcock:2024:GDF

[496] Gunnar Babcock and Daniel W. McShea. Goal directedness and the field concept. Philosophy of Science, 91(5):1435–1444, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/goal-directedness-and-the-field-concept/C0E700EA14F74F2948F6B280B4FC002E.

Sullivan:2024:DML

[497] Emily Sullivan. Do machine learning models represent their targets? Philosophy of Science, 91(5):1445–1455, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/do-machine-learning-models-represent-their-targets/
95A341566AF62EDDD5FAC2535DE59EC1.

Wilhelm:2024:PAE

[498] Isaac Wilhelm. The propositional account of effective theories. Philosophy of Science, 91(5):1456–1466, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.

cambridge.org/core/journals/philosophy-of-science/article/propositional-
account-of-effective-theories/3B605963A5151B33056D9A0E9161B42A.

Thompson:2024:PDM

[499] Morgan Thompson. Path-dependence in measurement: A problem for coherentism. Philosophy of Science, 91(5):1467–1477, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/pathdependence-in-measurement-
a-problem-for-coherentism/12067B2CBB1FA13B661EEBC7F0B201CA.

Bradley:2024:RRS

[500] Clara Bradley. The representational role of sophisticated theories. Philosophy of Science, 91(5):1478–1487, December 2024. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/representational-role-of-sophisticated-theories/
C0AAADDE9C8D79CC89FECD1BA125D71A.

Thornton:2024:EPD

[501] Suzanne Thornton and Minge Xie. An exploration of parameter duality in statistical inference. Philosophy of Science, 91(5):1488–1497, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/an-exploration-of-parameter-duality-in-statistical-
inference/D4DEB2C402A4F2DE217EE98BDAFEB941.

Mayo-Wilson:2024:BB

[502] Conor Mayo-Wilson. Bamboozled by Bonferroni. Philosophy of Science, 91(5):1498–1508, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/bamboozled-by-bonferroni/
0F570CAC41364F1F4D808A419A881277.

Anonymous:2024:PVIi

[503] Anonymous. PSA volume 91 issue 5 cover and front matter. Philosophy of Science, 91(5):f1–f6, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
5-cover-and-front-matter/A8DFD8D84E968E5681CB8E18D46F061A.

Anonymous:2024:PVIj

[504] Anonymous. PSA volume 91 issue 5 cover and back matter. Philosophy of Science, 91(5):b1–b5, December 2024. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-91-issue-
5-cover-and-back-matter/E80735DE0BEED6B019934DB4067D5A5F.

Politi:2025:CRS

[505] Vincenzo Politi. The collective responsibilities of science: Toward a normative framework. Philosophy of Science, 92(1):1–18, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/collective-responsibilities-of-science-toward-
a-normative-framework/8321A503C26656F3E7D51E4B7CCFA34F.

Fresco:2025:TSC

[506] Nir Fresco, Marc Artiga, and Marty J. Wolf. Teleofunction in the service of computational individuation. Philosophy of Science, 92(1): 19–39, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print),

1539-767X (electronic).  URL https://www.cambridge.org/core/
journals/philosophy-of-science/article/teleofunction-in-the-
service-of-computational-individuation/EF8F525BD640599468209C46EA0F8B72.

Kopecky:2025:IBA

[507] Felix Kopecky and Gregor Betz. Inconsistent belief aggregation in diverse and polarised groups. Philosophy of Science, 92(1):40–58, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/inconsistent-belief-aggregation-in-diverse-and-
polarised-groups/5B6A92D7317637E6F64CB78E9BE0D117.

Heine:2025:MCS

[508] Kyle B. Heine and Elay Shech. The mitonuclear compatibility species concept, intrinsic essentialism, and natural kinds. Philosophy of Science, 92(1):59–81, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/mitonuclear-
compatibility-species-concept-intrinsic-essentialism-and-natural-
kinds/7AF3D7B708461CE6892F5FCD85FE6157.

Coates:2025:DIH

[509] Matthew Coates. Does it harm science to suppress dissenting evidence? Philosophy of Science, 92(1):82–99, January 2025. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/does-it-harm-science-to-suppress-dissenting-
evidence/D24409D79DC77F123D33CBAB9CD9345C.

Koberinski:2025:BEF

[510] Adam Koberinski. The breakdown of effective field theory in particle physics: Lessons for understanding intertheoretic relations. Philosophy of Science, 92(1):100–120, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/breakdown-of-effective-
field-theory-in-particle-physics-lessons-for-understanding-intertheoretic-
relations/0AAA50D85D48873933E2740F1C4A045E.

Aronowitz:2025:LVS

[511] Sara Aronowitz. Locating values in the space of possibilities. Philosophy of Science, 92(1):121–140, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/locating-values-
in-the-space-of-possibilities/ED37B8FE6A9D47654AB0C88705D3910A.

Douven:2025:SLN

[512] Igor Douven. Social learning in neural agent-based models. Philosophy of Science, 92(1):141–161, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/social-learning-
in-neural-agentbased-models/88677994DD97E8824060FEF77943B641.

Kaiser:2025:ERI

[513] Marie I. Kaiser and Katie H. Morrow. On the epistemic roles of the individualized niche concept in ecology, behavioral biology, and evolutionary biology. Philosophy of Science, 92(1):162–184, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/on-the-epistemic-roles-of-the-
individualized-niche-concept-in-ecology-behavioral-biology-and-
evolutionary-biology/C7E1B4EF8FDA911BFEBD9ED0B6C0A9EB.

Barack:2025:CNF

[514] David L. Barack. Context and neural function. Philosophy of Science, 92(1):185–203, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.

org/core/journals/philosophy-of-science/article/context-and-
neural-function/3B0B93752A06F2B8F072FE6277D44F61.

Schindler:2025:ESS

[515] Samuel Schindler. Explanatory stabilization in situations of evidential uncertainty. Philosophy of Science, 92(1):204–222, January 2025.

CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/explanatory-stabilization-in-situations-of-evidential-
uncertainty/3AB708A38B6D3A3945FCF8FE045701D7.

Cifuentes:2025:FRB

[516] Arturo Cifuentes and Ventura Charlin. Further reflections based on the Black–Scholes (B-S) model. Philosophy of Science, 92(1):223–228, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/further-reflections-based-on-the-blackscholes-
bs-model/D9C3070AC13AA95A270D8DA9151C4E29.

Field:2025:RDH

[517] Grace Field. Review of Dominik Hangleiter, Jacques Carolan, and Karim P. Y. Thébault’s Analogue Quantum Simulation: A New Instrument for Scientific Understanding. Cham: Springer (2022), XI, 148 pp. £54.99 (hardback), £43.99 (ebook). Philosophy of Science, 92(1):229– 232, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-dominik-hangleiter-jacques-
carolan-and-karim-p-y-thebaults-analogue-quantum-simulation-
a-new-instrument-for-scientific-understanding-dominik-hangleiter-
jacques-carolan-and-karim-p-y-thebault-analogue-quantum-simulation-
a-new-instrument-for-scientific-understanding-cham-springer-
2022-xi-148-pp-5499-hardback-4399-ebook/A804A7BE7190C53D57F52B4E4CBA1D36.

Psillos:2025:RAH

[518] Stathis Psillos and Stavros Ioannidis. Review of Andreas Hüttemann’s A Minimal Metaphysics for Scientific Practice. Cambridge University Press (2021), viii+232 pp. $99.99 (hardcover). Philosophy of Science, 92(1): 233–236, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-

767X (electronic). URL https://www.cambridge.org/core/journals/
philosophy-of-science/article/review-of-andreas-huttemanns-a-
minimal-metaphysics-for-scientific-practice-andreas-huttemann-
a-minimal-metaphysics-for-scientific-practice-cambridge-university-
press-2021-viii232-pp-9999-hardcover/7F979DCF4C554390F4BB2E2417DA3251.

Anonymous:2025:RWH

[519] Anonymous. Review of Werner Heisenberg’s Reality and its Order — corrigendum. Philosophy of Science, 92(1):237, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/review-of-werner-heisenbergs-reality-and-its-
order-corrigendum/21410DA022A813360485CE8DC863C4DE. See [332].

Watkins:2025:PPB

[520] Aja Watkins. Paleoclimate proxies and the benefits of disunity — erratum. Philosophy of Science, 92(1):238, January 2025. CO-DEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic).

URL https://www.cambridge.org/core/journals/philosophy-of-
science/article/paleoclimate-proxies-and-the-benefits-of-disunity-
erratum/BE8D3AF8D61C48D60AEBF47ABDABDB9B. See [442].

Anonymous:2025:PVIa

[521] Anonymous. PSA volume 92 issue 1 cover and front matter. Philosophy of Science, 92(1):f1–f4, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-92-issue-
1-cover-and-front-matter/B75969610318AEAB349E2E4835D722DE.

Anonymous:2025:PVIb

[522] Anonymous. PSA volume 92 issue 1 cover and back matter. Philosophy of Science, 92(1):b1–b2, January 2025. CODEN PHSCA6. ISSN 0031-8248 (print), 1539-767X (electronic). URL https://www.cambridge.org/

core/journals/philosophy-of-science/article/psa-volume-92-issue-
1-cover-and-back-matter/753C1C18753A5EEC065C31B364BC3FB9.